1 JD Direct Impact Activites - Summary of Functions performed General observations - Issues/Comments Time series check (.7-2.0hz) (hf:7-9hz) avg. Wind speed (m/s) Significant signals (>.1pa) (340+/- 40m/s) Power spectra data check SOH check   I57US -> UCSD Daily Uptime Profile2 -> UCSD Daily Uptime I57US -> IDC Daily Uptime I57US -> NDC Daily Uptime   I57US -> UCSD Monthly Uptime Profile2 -> UCSD Monthly Uptime I57US -> IDC Monthly Uptime I57US -> NDC Monthly Uptime   I57US -> UCSD Annual Uptime Profile2 -> UCSD Annual Uptime I57US -> IDC Annual Uptime I57US -> NDC Annual Uptime   ID of report submitted/updated inventory changes Indirect Activites - Journal Direct Activities - Trip report
1/1/2014 1 - H4 BDA 20s spikes
- L2 ramp burst: 06:00, 11:30, 16:50, 17:30
- L4 hf ramp burst: 02:40
- L1 hf ramp burst: 04:15, 18:45
1.0928 - elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 4 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000   Fix scripts
- Modify all scripts including those for automated scp of reftek data and all soh processessing scripts for 2014
- Attempt to fix matlab license error
- Contact netops
1/2/2014 2 - H4 BDA 20s spikes
- L2 ramp burst: 08:40, 16:00, 16:45
- L4 hf ramp burst: 03:35
- L1 hf ramp burst: 05:30, 18:40
1.0999 - elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 3 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000   Daily log
- Uptimes: 357 – 365
- All: 357

December timesheet
- Fill out funding spreadsheet
- Send numbers to michael before filling out actual timesheet

Status meeting with michael
- Asked me to email ultisat regarding I57 UPS

Ultisat
- Sent email and asked for update
- Sent 2nd email to randy responding to him
1/3/2014 3 - H4 BDA 20s spikes
- L2 ramp burst: 07:20, 15:50, 16:45
- L4 hf ramp burst: 03:35
- L1 hf ramp burst: 05:25, 18:40
- H4 hf ramp burst: 13:30
1.008 - 21:44 (0.1pa 298) - elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 3 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000   MR - 99504 SOH script
- Error - restart

Reports
- Submit MR
1/4/2014 4 - H4 BDA 20s spikes
- L4 hf ramp burst: 04:30
1.5104 - elec noise:
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 2 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000  
1/5/2014 5 - H4 BDA 20s spikes
- L2 ramp burst: 09:25, 15:25, 17:10
- L4 hf ramp burst: 05:00
- L1 hf ramp burst: 07:20, 19:00
1.1046 - 05:37 (0.1pa 278)
- 23:13
(0.2pa 205)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 3 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000  
1/6/2014 6 - H4 BDA 20s spikes
- L2 ramp burst: 05:25, 10:10, 17:20, 18:35
- H3 ramp burst: 15:20, 17:50
- L4 hf ramp burst: 02:50
- L1 hf ramp burst: 04:00, 19:50
1.1337 - 20:28 (0.1pa 051)
- 22:52
(0.2pa 001)
- elec noise: H3, L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 3 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000   Soh script
- Error again, find problem and fix script
- Run for past 2 days
- Install mac updates and restart computer

Michael status call
- I’ll start work on summary report next week and give to michael for review before submission

Daily log
- All: 358 – 361
1/7/2014 7 - H4 BDA 20s spikes
- L2 ramp burst: 05:05, 09:10, 17:45, 18:35
- L4 hf ramp burst: 02:20
- L1 hf ramp burst: 03:40, 19:30
0.97737 - 00:29 (0.1pa 325)
- 03:52 (0.1pa 014)
- 18:35
(0.1pa 014)
- 18:44
(0.2pa 018)
- 23:29
(0.3pa 345)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all except H3 clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000  
1/8/2014 8 - H4 BDA 20s spikes
- L2 ramp burst: 10:25, 15:15, 16:15, 17:15
- L1 hf ramp burst: 05:50, 18:50
1.2079 - 21:04 (0.3pa 220)
- 22:32
(0.2pa 257)
- 23:35
(0.1pa 230)
- elec noise: L1-L3
- wnrs noise: L1,L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000  
1/9/2014 9 - H4 BDA 20s spikes
- L2 ramp burst: 06:00, 13:00, 17:15, 17:50
- L1 hf ramp burst: 19:20
1.0616 - 00:10 (0.2pa 028)
- 02:01 (0.3pa 217)
- 04:25
(1.0pa 213)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 3 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000   Lab
- Clean back bench and connect all required items for continual charging
- Take down thorpe and put in cabinent with parche (which was shut down)

Daily log
- All: 362 – 365

Soh script
- Modify to handle the full.wfdisc automatically
- Cleanup some of the script
1/10/2014 10 - H4 BDA 20s spikes
- L2 ramp burst: 06:45, 16:40
- L1 hf ramp burst: 04:15, 18:05
- L4 hf ramp burst: 20:00
1.154 - 02:00 (0.3pa 224) - elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all except H3 clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000   Soh script
- Troubleshoot, fix yesterdays updates

57 summary report
- Complete and send to michael for review
1/11/2014 11 - H4 BDA 20s spikes
- L2 ramp burst: 09:00, 16:30
- L4 hf ramp burst: 04:10, 20:15
- L1 hf ramp burst: 07:10, 18:20
.99472 - 01:37 (0.1pa 289)
- 20:11
(0.2pa 049)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000  
1/12/2014 12 - H4 BDA 20s spikes
- L4 hf ramp burst: 05:55
- L1 hf ramp burst: 15:00, 16:50
1.5415 - elec noise:
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 1 frame showing H1 GPS unlocked
- H3 BDF out of range
- H4 BDF out of range
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000  
1/13/2014 13 - H4 BDA 20s spikes
- L2 ramp burst: 08:50, 15:20, 16:55, 17:10
- L4 hf ramp burst: 02:50
- H3 hf ramp burst: 16:50
1.1077 - 00:08 (0.1pa 278)
- 17:48 (0.2pa 353)
- 18:19
(0.2pa 291)
- 19:05
(0.1pa 016)
- 19:19
(0.1pa 015)
- 21:36
(0.1pa 016)
- elec noise: H3, L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000   Daily log
- All: 1-5
1/14/2014 14 - H4 BDA 20s spikes
- L2 ramp burst: 07:10
- L1 hf ramp burst: 05:10
1.8977 - 00:20 (0.2pa 295)
- 01:02 (0.2pa 295)
- 04:14
(0.1pa 280)
- 17:23
(0.1pa 278)
- 17:59
(0.1pa 280)
- 19:40
(0.2pa 319)
- 20:21
(0.2pa 101)
- 20:51
(0.2pa 197)
- 21:00
(0.4pa 202)
- elec noise: L1, L2
- wnrs noise: H2, L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000   Daily log
- All: 6-8

Michael Status
- send update emails on all
1/15/2014 15 - H4 BDA 20s spikes
- L2 ramp burst: 09:55, 16:30
- L1 hf ramp burst: 07:30, 18:45
1.1719 - 00:56 (0.2pa 326)
- 01:37 (0.1pa 279)
- 15:38
(0.1pa 119)
- 18:13
(0.3pa 291)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000  
1/16/2014 16 - H4 BDA 20s spikes
- L4 hf ramp burst: 03:40
- L1 hf ramp burst: 06:10
1.3806 - 00:51 (0.1pa 292)
- 00:59
(0.2pa 290)
- 01:16
(0.1pa 325)
- 01:41
(0.2pa 014)
- 01:53
(0.2pa 014)
- 03:18
(0.2pa 107)
- elec noise: L1, L2
- wnrs noise: H2, L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 16 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000   Daily log
- All: 9-10
1/17/2014 17 - H4 BDA 20s spikes
- L2 ramp burst: 12:50, 16:25
- L4 hf ramp burst: 03:50
- L1 hf ramp burst: 08:00, 19:00
1.1881 - 00:52 (0.2pa 296)
- 02:32 (0.1pa 025)
- 17:20
(0.1pa 295)
- 17:46
(0.2pa 292)
- 18:13
(0.4pa 293)
- 20:45
(0.1pa 315)
- 21:07
(0.2pa 301)
- 21:32
(0.3pa 334)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000   Daily log
- all: 11-16
1/18/2014 18 - H4 BDA 20s spikes
- L2 ramp burst: 09:45, 16:40
- L4 hf ramp burst: 03:55
- L1 hf ramp burst: 04:55, 19:35
1.047 - 00:45 (0.1pa 297)
- 01:03 (0.5pa 221)
- 01:15
(0.3pa 292)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000  
1/19/2014 19 - H4 BDA 20s spikes
- L2 ramp burst: 09:40, 17:20
- L4 hf ramp burst: 03:55
- L1 hf ramp burst: 04:15, 20:10
.98239 - elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000  
1/20/2014 20 - H4 BDA 20s spikes
- L2 ramp burst: 07:10, 16:20, 17:05
- L4 hf ramp burst: 02:30
- L1 hf ramp burst: 03:10, 19:50
1.0259 - 01:14 (0.1pa 326)
- 16:28
(0.2pa 277)
- 18:00
(0.2pa 220)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000   Soh script
- Could not fix from home so had to come in
- Found that umoya nfs server was not working
- Restarted umoya nfs and then restarted hal computer and reconnected to all servers
- Waited and confirmed all working
1/21/2014 21 - H4 BDA 20s spikes
- L2 ramp burst: 09:40, 16:50
- L4 hf ramp burst: 03:50, 22:40
- L1 hf ramp burst: 04:35, 20:45
.99408 - 05:33 (0.1pa 278)
- 17:40
(0.3pa 297)
- 21:12
(0.1pa 299)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 2 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000   Daily log
- All: 16-18

Express Card
- Look up purchase that is being questioned and give info to lenore
1/22/2014 22 - H4 BDA 20s spikes
- L2 ramp burst: 12:30, 16:15
- L4 hf ramp burst: 01:40
- L1 hf ramp burst: 02:40
.98411 - 00:25 (0.1pa 278)
- 00:54 (0.2pa 298)
- 16:05
(0.3pa 274)
- 17:37
(0.1pa 293)
- 19:28
(0.2pa 277)
- 21:43
(0.2pa 278)
- 22:24
(0.8pa 267)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all except L3 clk diff to large
- 3 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000   Daily log
- All: 19-21

Office computers
- Install second UPS
1/23/2014 23 - H4 BDA 20s spikes
- L2 ramp burst: 09:00, 16:40
- L1 hf ramp burst: 05:10, 19:40
1.23 - 02:11 (0.1pa 274)
- 04:09
(0.1pa 275)
- 07:48
(0.1pa 277)
- 08:10
(0.2pa 279)
- 17:40
(0.2pa 296)
- 18:07
(0.2pa 266)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000   Daily log
- All: 22
1/24/2014 24 - H4 BDA 20s spikes
- L2 ramp burst: 10:40, 17:10
- L1 hf ramp burst: 04:00, 21:00
1.0978 - 21:40 (0.1pa 336)
- 22:22
(0.2pa 340)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000   Daily log
- All: 23
1/25/2014 25 - 17:40 - purge frames restart computer - Forwarder outage: 11:11 - 17:40 - H4 BDA 20s spikes
- L2 single spike: 19:05
- L1 hf ramp burst: 02:20
- L2 hf ramp burst: 15:30
.97193 - elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 1277 frames from I57 to UCSD missing
- 1277 frames from I57 to NDC missing
- 1277 frames from I57 to IDC missing
  72.752 100.000 73.032 72.752 x 98.910 100.000 98.921 98.910 x 98.910 100.000 98.921 98.910   I57US outage
- I57 backup computer stopped sending data at 11:15
- logs show forwarder not enabled and kept trying to find a data record from 10/18/13 6:57
- continued to initialize forwarder for next 6 hours and then getting stuck on data record mentioned above
- could not vnc or ping the computer
- at 17:10 used baytech to switch from backup to primary computer
- still could not vnc or ping, no indication that primary computer is able to connect to umoy
- all above attempts to vnc or ping were either from home computer or hal2008. When arrived at work I COULD vnc from devi.
- 18:02 - Primary computer is sending data but reporting sql exception and continues to restart L4 thread
- 18:04 – stop all on primary computer
- 18:05 – switch back over to backup computer -> Still stuck on bad data record file mentioned above
- 18:17 – stop all, clear queue/fwd files and restart -> Working
1/26/2014 26 - H4 BDA 20s spikes
- L2 ramp burst: 09:00, 16:15
- L4 hf ramp burst: 04:10
- L1 hf ramp burst: 05:00, 18:40
1.0072 - 15:53 (0.1pa 207) - elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 98.952 100.000 98.963 98.952 x 98.952 100.000 98.963 98.952  
1/27/2014 27 - H4 BDA 20s spikes
- L2 ramp burst: 14:25, 16:00
- L1 hf ramp burst: 08:50
1.1857 - 05:48 (0.1pa 280)
- 17:52
(0.2pa 353)
- 18:04
(0.1pa 033)
- 18:07
(0.2pa 265)
- 18:09
(1.0pa 265)
- 18:15
(0.2pa 355)
- 18:27
(1.8pa 266)
- 18:37
(0.2pa 261)
- 19:19
(0.2pa 332)
- 19:27
(0.1pa 352)
- 21:20
(0.8pa 257)
- 22:11
(0.4pa 255)
- 22:49
(0.4pa 259)
- elec noise: L2, L4
- wnrs noise: L1, L3
- 8hz bump: L1-L4
- all except H2 clk diff too large   100.000 100.000 100.000 100.000 x 98.991 100.000 99.001 98.991 x 98.991 100.000 99.001 98.991   PR - 100077 Daily log
- All: 24-26

I57 outage
- Submit PR update

Secure keys
- Discuss process with randy on phone
- May go out this week to generate new keys

Michael status call
- I57 outage
- Michael would like the script to continue to notify him and notify when problem fixed.
- Vacation
- I57 key trip
1/28/2014 28 - H4 BDA 20s spikes
- L2 ramp burst: 11:55, 16:10
- L4 hf ramp burst: 05:30
- L1 hf ramp burst: 05:50
1.1429 - 04:50 (3.0pa 213)
- 16:37
(0.2pa 355)
- 16:55
(0.1pa 355)
- 17:04
(0.6pa 355)
- 18:28
(0.2pa 323)
- 18:37
(0.2pa 292)
- 21:27
(0.3pa 277)
- 21:57
(0.3pa 281)
- 22:14
(0.2pa 282)
- 22:50
(0.3pa 286)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 2 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 99.027 100.000 99.037 99.027 x 99.027 100.000 99.037 99.027   Daily log
- All: 27
1/29/2014 29 - H4 BDA 20s spikes
- L2 ramp burst: 13:40, 15:40
- L4 hf ramp burst: 06:20
- L1 hf ramp burst: 06:55
1.0802 - 00:31 (0.2pa 274)
- 16:38 (0.1pa 356)
- 19:34
(0.1pa 273)
- 20:04
(0.2pa 272)
- 20:14
(0.2pa 272)
- 20:25
(0.6pa 347)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 3 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 99.060 100.000 99.070 99.060 x 99.060 100.000 99.070 99.060   Daily log
- All: 28

57 outage
- Download and compress log. Send to gary with explanation and ask about modifying the watchdog to purge files

Pinon prep
- Build cables
- Gather equipment
- Fuel truck
1/30/2014 30 - H4 BDA 20s spikes 1.9089 - 00:05 (0.4pa 331)
- 00:15 (0.2pa 332)
- elec noise: L2
- wnrs noise: L1-L3, H1
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.092 100.000 99.101 99.092 x 99.092 100.000 99.101 99.092  
1/31/2014 31 - Profile2 windows updates
- Profile2 free disk space
- Profile2 dbconfig changes
- H2 add vba channel
- L2 add vba channel and cable
- L2 digitizer damaged, replace with I57CF digitizer
- I57CF channels offline - H4 BDA 20s spikes - N/A - all clk diff too large
- H1 BDF out of range
- H4 BDF out of range
- 8 frames from I57 to IDC missing
- 7 frames from I57 to NDC missing
- 8 frames from I57 to UCSD missing
- 22 frames from Profile2 to UCSD missing
  99.035 98.874 99.815 99.056 x 99.090 99.964 99.124 99.091 x 99.090 99.964 99.124 99.091   UPS
- After the reboot that was performed on the 25th the USB port was refreshed and connectivity to the UPS was restored
- A self check states that the battery must be replaced
- Runtime check claims it still has 41min of runtime with current load
- The check might be based on age and not any actual physical parameters. A 41min runtime is more than enough

Profile2
- Installed 11 windows updates
- Reconfigured smart24db to allow 2 aux channels on H2 and L2
- I/O aux – VBA
- Cleared fwdr queues and restart computer
- After restart smartserver was complaining about IO error for L2 channel:
- Stopall
- Repair tables
- Restart
- Working as of 16:16
- Computer showing disk space errors:
- only 300mb of free memory left
- cleared all forwarder logs before 2014
- now has 6gb free

H2
- use smart24config to reconfigure digitizer to add IO aux VBA channel

L2
- use smart24config to reconfigure digitizer to add IO aux VBA channel
- Upon entry to vault notice that digitizer LCD screen reporting that digitizer is stuck in bootloader mode:
- Cannot connect with smart24config from crf
- Cannot ping from crf
- Several power cycles had no effect
- Attempt power cycle with on/off buttons but no effect
- Attempt reboot with on/off/user buttons but no effect
- Unable to scroll through soh screens on lcd display although the manual says you should be able to in boot loader mode
- Reconfigured I57CF digitizer (1222) as L2.
- Swapped damaged L2 digitizer (1229) with I57CF digitizer (1222) – works fine
- 17:07 – finish installing new VBA cable. Solder in cable with in-line voltage divider to the existing IO2 connector
                    x x x x x x x x x x x x x x          
2/1/2014 32 - I57CF channels offline - H4 BDA 20s spikes
- L2 ramp burst: 04:10, 20:15, 21:25
1.5994 - 18:33 (0.6pa 315)
- 19:19
(1.5pa 315)
- elec noise: L2
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.118 99.965 99.151 99.119  
2/2/2014 33 - I57CF channels offline - H4 BDA 20s spikes
- L2 ramp burst: 02:05, 04:40, 20:30
- H3 ramp burst: 16:40, 18:00
- L1 hf ramp burst: 00:30
1.1468 - 01:27(0.2pa 216)
- 19:28
(0.3pa 191)
- elec noise: L1, L2, H3
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.145 99.966 99.177 99.146   Soh Script
- Edit and run several times to compensate for I57CF being offline
2/3/2014 34 - I57CF channels offline - H4 BDA 20s spikes
- L2 ramp burst: 20:10
- L2 hf ramp burst: 01:15, 04:30
1.3921 - 16:31 (0.7pa 255) - elec noise: L2
- wnrs noise: L1-L3, H1
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.170 99.967 99.201 99.171   MR - 100201 Daily log
- Reprocess 32 - 33

Soh script
- Continue to fix to compensate for I57CF being offline

Timesheet
- Fillout, confirm with michael, copy and submit

Smart forwarder error
- Email gary and tell him I was out and that I will call later this morning
- Call gary and leave message

I57 field work
- Write trip report
- Update daily log with actions
- Submit PR
- Fwd to CMR with additional comments

Reports
- Submit January MR

Michael status call
- Discuss all

Umoya ntp security
- Research online
- Edit ntp.conf file to restrict ntp access to the ucsd time server only and ignore all others:
-Discovered that yast cannot do this for some reason via gui and when I load it after doing it manually it erases changes. For this reason I cannot use yast again in the future when working the ntp configuration.
- Email wayne with changes and ask if enough
- Work with say-ho to further modify ntp.conf file and then restart ntp service
2/4/2014 35 - I57CF channels offline - H4 BDA 20s spikes
- L2 ramp burst: 00:55, 04:00, 20:35, 21:25
- H3 hf ramp burst: 16:40
1.7349 - 00:13 (0.5pa 270)
- 00:39 (0.2pa 273)
- 19:56 (0.2pa 197)
- elec noise: L2, H3
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- H3 BDF out of range
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.194 99.968 99.224 99.195   PR - 100077
PR - 100200
Filel2a
- Server went down last night causing soh script to fail
- Contact netops for help
- Hardware failure last night caused all the problems
- Brought back online by netops at 11am and restarted all
- Rerun soh script for yesterday

L2 digitizer outage
- Submit ccn
- Build smart24 pwr cable
- Still boots up in bootloader mode
- Attempt serial connection but could not connect
- Troubleshoot with gary determine it was a short in the touchpad
- Update inventory, and digitizer notes
- Update PR
- Place order for new touchpad

Container
- Took load of equipment including MB cases and erdc enclosures to the container
- Recorded in inventory
2/5/2014 36 - I57CF channels offline - H4 BDA 20s spikes
- L2 ramp burst: 01:15, 05:20, 19:30, 20:10
- H3 hf ramp burst: 16:50
.97459 - 00:21 (0.1pa 328)
- 00:54 (0.2pa 277)
- 02:12
(0.1pa 242)
- elec noise: H3, L2
- wnrs noise: L1-L2
- 8hz bump: L1-L4
- all except H2 clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.216 99.969 99.246 99.217  
2/6/2014 37 - H2 outage:
-
I57CF channels offline
- H4 BDA 20s spikes
- L2 ramp burst: 02:55, 06:50
1.2993 - elec noise: L2
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all except L3 clk diff too large   96.620 100.000 100.000 96.608 x 99.437 100.000 100.000 99.435 x 99.146 99.970 99.266 99.146  
2/7/2014 38 - H2 outage:
-
I57CF channels offline
- H4 BDA 20s spikes 2.8764 - elec noise: L2
- wnrs noise: L1-L3, H1-H3
- 8hz bump: L1-L4
- all clk diff too large
- L1, L3, H1, H2, H4 BDF out of range
- 5 frames from I57 to IDC delayed
- 5 frames from I57 to NDC delayed
- 5 frames from I57 to UCSD delayed
  92.863 100.000 100.000 92.863 x 98.498 100.000 100.000 98.496 x 98.981 99.970 99.285 98.981  
2/8/2014 39 - I57CF channels offline - H4 BDA 20s spikes
- L2 ramp burst: 01:45, 17:40, 18:10
1.8452 - elec noise: L2
- wnrs noise: L1-L3, H1-H3
- 8hz bump: L1-L4
- all except L2 clk diff too large
- 4 frames from I57 to IDC delayed
  100.000 100.000 100.000 100.000 x 98.685 100.000 100.000 98.684 x 99.007 99.971 99.304 99.007  
2/9/2014 40 - I57CF channels offline - H4 BDA 20s spikes
- L2 ramp burst: 10:30
- L2 hf ramp burst: 16:00
2.1126 - elec noise: L2
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 98.832 100.000 100.000 98.830 x 99.032 99.972 99.321 99.032  
2/10/2014 41 - I57CF channels offline - H4 BDA 20s spikes 2.0361 - elec noise: L2
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- H1 BDF out of range
  100.000 100.000 100.000 100.000 x 98.948 100.000 100.000 98.947 x 99.055 99.973 99.338 99.056   Soh Script
- repair and start re-processing last 5 days
2/11/2014 42 - I57CF channels offline - H4 BDA 20s spikes
- L2 ramp burst: 13:00, 16:10
- L4 hf ramp burst: 09:20
- L1 hf ramp burst: 11:40
1.2012 - elec noise: L2
- wnrs noise: L1-L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.044 100.000 100.000 99.043 x 99.078 99.973 99.354 99.078  
2/12/2014 43 - 17:11 - restart umoya - I57CF channels offline - H4 BDA 20s spikes
- L2 ramp burst: 10:50, 16:05
- L4 hf ramp burst: 07:10
- L1 hf ramp burst: 08:50
1.0189 - elec noise: L1-L2
- wnrs noise: L1-L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.124 100.000 100.000 99.123 x 99.099 99.974 99.369 99.100   Daily log
- All: 29-30

Soh script
- Repair and reprocess last 2 days several times to get past errors

Smart24 damage
- Receive new touchpad – submit purchase paperwork
- Remove old touchpad and install new one (new version!)
- Test – working well

Umoya nfs
- Restart computer
- Respond to wayne
- As108 & profile2 did not reconnect:
- Disable firewall -> They connected and backfilled
- Made additional changes to ntp.conf file, restarted – spoke with say-ho and confirmed working

Prep for pinon trip
- Fuel truck
- Load equip
2/13/2014 44 - Profile2 dbconfig changes
- I57 & Prof2 table repair and restart
- H2 vdiv configure & install
- L2 digitizer swap
- I57CF digitizer install
- H4 BDA 20s spikes
- L4 hf ramp burst: 14:30, 16:20
- N/A - all clk diff too large
- visit related flags
- 7 frames from I57 to IDC, UCSD, NDC missing
- 17 frames from profile2 to UCSD missing
  99.350 99.722 99.815 99.722 x 99.141 99.979 99.986 99.169 x 99.105 99.968 99.379 99.114   L2
- 18:30 - configure digitizer currently at L2 (1222) as I57CF over radio link. remove and install the repaired 1229 digitizer at L2
- 18:45 - install 1222 back at CRF and confirm via viewer and both servers that both digitizers are sending data

Profile2
- 18:55 - discover that profile2 computer had not been syncing clock and was 5min off, adjusted settings to look for our server (172.23.34.87) which worked and it synced. caused several port conflicts so had to stop all, clean tables and restart

All
- 19:30 - inspect L4, H1, H4, L1 interiors: L3 blocked

H2
- 20:00 - vdiv at H2: found that already connected and working to mass pos #1
- 20:20 - measurement shows voltage at 13.93 and being divided to 6.88 to mass pos #1
- 20:30 - deactivate io aux channel on h2 and activate mass pos #1
- 20:35 - io exceptions on both computers. stopall on both, repair tables on both, update dbconfig on profile2 for switch from io aux to mass position and then restart both computers
2/14/2014 45 - H4 BDA 20s spikes 1.2599 - 00:10 (0.5pa 207)
- 04:11 (0.1pa 274)
- 21:04
(0.5pa 207)
- elec noise:
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 42 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 99.202 99.980 99.987 99.228 x 99.125 99.969 99.392 99.133   CCN - 100236
PR - 100200
Daily log
- All: 31-35

Soh script
- Update to reflect soh channels are backonline
- Reprocess yesterday

Field work
- Write trip report
- Update log
- Update inventory
- Submit PR and CCN updates

L3 Access
- Ask Al if I could get access across stonecrest through new anza elect steel rope
- Al’s email doesn’t work, ask frank if he has new contact
2/15/2014 46 - H4 BDA 20s spikes
- L2 hf ramp burst: 02:05
1.2011 - elec noise: L2
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 4 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 99.256 99.981 99.988 99.280 x 99.144 99.969 99.406 99.152  
2/16/2014 47 - H4 BDA 20s spikes 1.2952 - elec noise: L2
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 3 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 99.302 99.983 99.988 99.325 x 99.162 99.970 99.418 99.170  
2/17/2014 48 - H4 BDA 20s spikes
- L4 hf ramp burst: 14:10
- L1 hf ramp burst: 16:10
1.2891 - elec noise: L1, L4
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 3 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 99.343 99.984 99.989 99.364 x 99.180 99.971 99.430 99.188  
2/18/2014 49 - H4 BDA 20s spikes
- L2 ramp burst: 14:50, 15:15
- L1 hf ramp burst: 16:30
1.5207 - elec noise: L2
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 9 frames from I57 to UCSD delayed
- 11 frames from profile2 to UCSD delayed
- 1 frame from profile2 to UCSD duplicated
- 5 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 99.380 99.985 99.990 99.400 x 99.196 99.971 99.442 99.204   CCN - 100236 Daily log
- All: 36 - 48

Ccn
- Respond to dina with sn of auth card in digitizer – confirm by logging into digitizer

Michael meeting
- Discuss status of all
- Send list of signals tracked

Ucr land
- Find Al contact info online and send another email
2/19/2014 50 - H4 BDA 20s spikes 1.3381 - 00:29 (0.1pa 257)
- 02:08 (0.3pa 293)
- 03:10
(0.2pa 012)
- 20:30
(1.2pa 243)
- 23:41
(2.0pa 232)
- 23:53
(0.5pa 012)
- elec noise:
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all except L3 clk diff too large
- 5 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 99.412 99.985 99.990 99.431 x 99.212 99.972 99.453 99.220   Daily log
- All: 49

Files
- Cleanup ndula, devi, flash drives
- Move all infrasound files to freeagent for constant use
- Zip and backup to leave on devit

Soh script
- Cleanup
- Start work on year processing
2/20/2014 51 - H4 BDA 20s spikes
- L2 ramp burst: 13:50, 15:55
- L4 hf ramp burst: 12:00
- L1 hf ramp burst: 13:00, 17:00
1.3484 - 04:16 (0.1pa 271)
- 06:35
(0.1pa 002)
- 18:51
(0.5pa 025)
- elec noise: L1-L2
- wnrs noise: L1-L2
- 8hz bump: L1-L4
- all clk diff too large
- 5 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 99.442 99.986 99.991 99.460 x 99.228 99.972 99.464 99.235   Daily log
- All: 50

Soh script
- Continue update work
- Finish “fullyear_batch” script which found max/min/avg for all channels for the past 50 days and put them in a matrix
- Finish “fullyear” script which automatically runs at end of each day to add that days max/min/avg for all channels to the fullyear matrix and then to create output plots in the year directory

Gci ups
- Ask randy which reports I need to submit
- Give them the correct shipping address

Vdiv channels
- Ask randy if OR is necessary
2/21/2014 52 - H4 BDA 20s spikes
- L2 ramp burst: 09:15, 12:55
- L4 hf ramp burst: 06:50, 18:30
- L1 hf ramp burst: 07:30, 16:55
- H4 hf ramp burst: 13:00
1.1374 - 18:22 (0.2pa 330) - elec noise: L1-L2
- wnrs noise: L1-L2
- 8hz bump: L1-L4
- all clk diff too large
-
720 frames from I57 to NDC delayed
- 5 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 99.468 99.987 99.991 99.485 x 99.243 99.973 99.474 99.250   Daily log
- All: 51

Hpwren outage
- Started at 11:49ut
- Ndc went offline at 12:30UT

Soh script
- Update to copy year trends to website

Website
- Update with links to year trends
- Fix grid lines on 2014 log
- Email michael with updates
2/22/2014 53 - H4 BDA 20s spikes
- L2 ramp burst: 10:35, 16:10
- L4 hf ramp burst: 07:20
- L1 hf ramp burst: 08:25, 17:35
1.0973 - elec noise: L1-L2
- wnrs noise: L1-L2
- 8hz bump: L1-L4
- all except H3 clk diff too large
- 4 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 99.492 99.987 99.992 99.509 x 99.257 99.974 99.484 99.264  
2/23/2014 54 - H4 BDA 20s spikes
- L2 ramp burst: 10:40, 16:15
- L4 hf ramp burst: 06:20
- L1 hf ramp burst: 08:10, 17:50
- H4 hf ramp burst: 10:40
1.1582 - 08:20 (0.1pa 284) - elec noise: L1-L2
- wnrs noise: L1-L2
- 8hz bump: L1-L4
- all clk diff too large
- 4 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 99.515 99.988 99.992 99.530 x 99.271 99.974 99.494 99.278  
2/24/2014 55 - H4 BDA 20s spikes
- L2 ramp burst: 10:30, 16:05
- L4 hf ramp burst: 06:45
- L1 hf ramp burst: 07:45, 17:45
- H1 hf ramp burst: 10:10
1.1589 - 22:56 (0.3pa 264)
- 23:21
(0.1pa 348)
- elec noise: L1-L2
- wnrs noise: L1-L2
- 8hz bump: L1-L4
- all clk diff too large
- 4 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 99.535 99.988 99.992 99.550 x 99.284 99.974 99.503 99.291   Daily log
- All: 52 – 54

Status meeting
- Email michael week schedule

Randy
- Email re: new gci ups
- Email re: voltage monitoring circuits
- Email re: moving old as108 equip out

Vault lids
- Start work on design for removable hinges
- Research hinges
- Order hinges from mcmaster

Pipe filter
- Redesign with simpler 10’ pipe lengths (since home depot sells at this length)
- Check material cost - $85
- Ask michael for permission to build in a couple of weeks

Soh script
- Continue work on plot/table/settings buttons
2/25/2014 56 - H4 BDA 20s spikes
- L2 ramp burst: 13:20, 16:05
- L4 hf ramp burst: 07:10
- L1 hf ramp burst: 08:20, 17:35
1.1298 - 18:07 (0.2pa 296)
- 18:43
(0.1pa 204)
- 19:06
(0.2pa 204)
- 19:40
(0.2pa 339)
- 20:05
(0.3pa 265)
- 20:26
(0.1pa 256)
- 21:03
(0.6pa 344)
- 22:22
(0.2pa 299)
- elec noise: L1-L2
- wnrs noise: L1-L2
- 8hz bump: L1-L4
- all clk diff too large
- 4 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 99.553 99.989 99.993 99.568 x 99.297 99.975 99.512 99.304   Daily log
- All: 55

Pipe filter
- Continue work on square design

Soh script
- Continue work on buttons
2/26/2014 57 - H4 BDA 20s spikes
- L2 hf ramp burst: 15:00
1.4296 - 01:22 (0.2pa 204)
- 01:37
(0.1pa 338)
- 19:58
(0.2pa 338)
- elec noise: L1-L2
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 3 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 99.571 99.989 99.993 99.584 x 99.309 99.975 99.520 99.316   Daily log
- All: 056

Soh script
- Continue work on buttons
2/27/2014 58 - H4 BDA 20s spikes 2.5921 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 3 frames showing H1 GPS unlocked
- H1 BDF out of range
- H2 BDF out of range
- H3 BDF out of range
- H4 BDF out of range
  100.000 100.000 100.000 100.000 x 99.586 99.990 99.993 99.600 x 99.321 99.976 99.529 99.328   CCN - 100236
OR - 100837
Daily log
- All: 057

Wnrs
- Email michael with additional figures and more explanation of the square design test

Authentication cards
- Check all digitizers and reconcile database
- Respond to ccn

Gci ups
- Receiver equipment
- Notify ultisat
- Submit OR
2/28/2014 59 - H4 BDA 20s spikes
- L1 single spike: 10:20
2.9946 - 00:31 (0.2pa 299)
- 01:16 (0.1pa 272)
- elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 3 frames showing H1 GPS unlocked
- L2 BDF out of range
- L3 BDF out of range
- L4 BDF out of range
- H1 BDF out of range
- H2 BDF out of range
- H3 BDF out of range
- H4 BDF out of range
  100.000 100.000 100.000 100.000 x 99.601 99.990 99.993 99.614 x 99.333 99.976 99.537 99.339   Archive outage
- Attempt to bring matlab scripts online
- Communicate with netops
- Work on system remotely but cannot resolve
                    x x x x x x x x x x x x x x          
3/1/2014 60 - Power surge at 00:30 caused AC unit to stop functioning - H4 BDA 20s spikes 2.6806 - elec noise: L2
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- H1 BDF out of range
- H2 BDF out of range
- H3 BDF out of range
- H4 BDF out of range
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.344 99.977 99.544 99.350  
3/2/2014 61 - AC unit not working - H4 BDA 20s spikes
- L2 ramp burst: 04:05, 19:25, 20:10
- L1 hf ramp burst: 20:00
1.6618 - elec noise: L2
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.354 99.977 99.552 99.361  
3/3/2014 62 - AC unit not working - H4 BDA 20s spikes
- L2 ramp burst: 02:00, 07:00, 18:10, 19:00
- L1 hf ramp burst: 03:10, 18:50
- H4 hf ramp burst: 11:40
1.1891 - 21:23 (0.6pa 222) - elec noise: H1, L1-L3
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.365 99.977 99.559 99.371   Archive connection
- Can use terminal to connect but nothing else (finder, matlab, nfs connection, etc…)
- Research, try restarts, etc…
- Discuss with say-ho on phone

Status meeting
- Meet with michael, discuss all

Vandenberg
- Send michael latest report on upcoming launches

Timesheet
- Ask michael for confirmation
- Fill out, copy and submit
3/4/2014 63 - AC unit not working - H4 BDA 20s spikes
- L2 ramp burst: 06:30, 16:45
- L4 hf ramp burst: 03:00
- L1 hf ramp burst:06:40
1.0657 - 17:14 (0.3pa 358)
- 17:49
(0.2pa 350)
- 18:19
(0.2pa 283)
- 19:11
(0.1pa 119)
- 23:14
(0.1pa 004)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.375 99.978 99.566 99.381   Gci UPS
- Respond to hamid with scheduled install and several follow-up emails

Clean lab
- Organize software, etc…

Wnrs
- Purchase pvc materials at home depot, unload and submit receipts

Archive outage:
- Happens because of my older mac osx version
- Work with say-ho on temporary fix
- Company is working on update which should be avail next week

Re-process:
- Modify script for new function inputs and run for last 5 days
3/5/2014 64 - Don reset AC breaker at 23:05 - AC unit not working - H4 BDA 20s spikes
- L2 ramp burst: 12:00, 15:45
- L4 hf ramp burst: 06:20
- H4 hf ramp burst: 09:10
- L1 hf ramp burst: 10:40
1.1447 - 02:17 (0.3pa 221)
- 03:05
(0.1pa 354)
- 16:56
(0.1pa 345)
- 19:43
(0.1pa 296)
- 23:37
(1.7pa 266)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- visit related flags
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.385 99.978 99.573 99.391   Archive
- One of my scripts last night created a large all.wfdisc file on day 58 that filled the archive
- Say-ho gave me extra space so I can find and delete
- Then re-process day 62, 63
- Fix full year and re-process to give accurate plots

Air conditioner
- Went offline during the storm, the crf is slowly heating up
- Left messages at all 3 numbers for don
3/6/2014 65 - H4 BDA 20s spikes 1.3422 - 08:30 (0.1pa 266)
- 08:41
(0.1pa 265)
- 16:34
(0.1pa 284)
- 19:59
(0.1pa 297)
- elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all except L3, L4 clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.394 99.978 99.579 99.400   MR - 101024 Archive
- Nfs fix did not work
- Password problem stopped me from logging back in to afp, resolve with netops
- Reprocess 64

Air conditioner
- Don fixed breaker at 3:05 local yesterday

Daily log
- All: 58-61

MR
- Organize emails
- Submit report
3/7/2014 66 - H4 BDA 20s spikes 2.0091 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 751 frames from I57 to NDC missing
- 49 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.403 99.979 99.586 99.409   Archive
- restart soh script after recovery from network outage
3/8/2014 67 - H4 BDA 20s spikes 1.8046 - 07:16 (0.1pa 280) - elec noise
- wnrs noise: H1-H2, L1-L2
- 8hz bump: L1-L4
- all clk diff too large
- 4320 frames from I57 to NDC missing
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.412 99.979 99.592 99.418  
3/9/2014 68 - H4 BDA 20s spikes
- L4 hf ramp burst: 10:35
1.6251 - elec noise:
- wnrs noise: H1-H2, L1-L3
- 8hz bump: L1-L4
- all except H3 clk diff too large
- 4320 frames from I57 to NDC missing
- 2 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.421 99.979 99.598 99.426  
3/10/2014 69 - H4 BDA 20s spikes
- L4 hf ramp burst: 07:40,
- L1 hf ramp burst: 14:40
1.2947 - elec noise: L1, L2
- wnrs noise: H1, L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 11760 frames from I57 to NDC delayed
- 9391 frames from I57 to NDC backfilled
- 2 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.429 99.980 99.604 99.435  
3/11/2014 70 - Install GCI UPS
- Add VBA & SQU channels
- calib values set to 0 - H4 BDA 20s spikes
- L4 hf ramp burst: 11:25, 17:40
1.5199 - 21:46 (0.2pa 256)
- 22:32 (
0.2pa 257)
- 23:31
(0.4pa 255)
- 23:55
(0.3pa 268)
- elec noise:
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 82 frames from I57 to IDC missing
- 82 frames from I57 to NDC missing
- 82 frames from I57 to UCSD missing
- 78 frames from profile2 to UCSD missing
- site visit related flags
  97.990 98.134 97.864 97.990 x 99.817 99.830 99.806 99.817 x 99.409 99.953 99.579 99.414   GCI
- 15:45 - 16:30 = swap UPS units
- 16:30 - 16:35 = connect gci equipment and network cable to new UPS
- 16:35 - 16:45 = accidentaly disconnect my UPS cable in the process above which killed power to crf equipment. forced to repair tables and restart software several times to restore communications
- 16:45 = both computers sending to all consumers without problems
- 18:40 = long call from ultisat (jim lang) about all the stuff they want me to do. told them I would have to come back next week.
- 18:54 = plug in backup gci modem to power and fe9

VBA
- 17:05 - 17:25 = stop all, add VBA channel to all digitizers, move from io to mp on L2, add squ on H2, update smartdb on profile2, startall
3/12/2014 71 - Large power draw around 18:00, possibly due to GCI UPS test being ran by Ultisat?
- calib values set to 0
- H4 BDA 20s spikes
- L4 hf ramp burst: 06:30
- L1 hf ramp burst: 12:50
1.5425 - 16:24 (0.2pa 269)
- 16:59
(0.3pa 268)
- 18:38
(0.5pa 319)
- 22:01
(0.3pa 347)
- 22:28
(0.3pa 260)
- 23:18
(0.2pa 259)
- 23:27
(0.2pa 327)
- elec noise: H2, L1-L2
- wnrs noise: L1-L2
- 8hz bump:L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.832 99.844 99.822 99.832 x 99.417 99.954 99.585 99.422   Token
- Fill out paperwork to register for my own token
- Scan and send in

Ctbt web portal
- Attempt to create new account – told I already have one
- Attempt to login – don’t have password of file
- Request username and receive
- Request for password failed
- Locked account after several login attempts
- Request assistance from ctbto
- Inform randy of all above

Ultisat
- Long email to michael/randy explaining the additional tasks they are requesting and if I should proceed

Trip return:
- Write trip report
- Unload truck

Daily log
- Re-process: 67, 68

Soh script
- Update to account for no data to one of the consumers and then re-run process above

Pinon smtp
- Place order for another year of service and send paperwork to lenore

Randy
- Update PR
3/13/2014 72 - calib values set to 0 - H4 BDA 20s spikes
- L2 ramp burst: 11:55, 14:55
- L4 hf ramp burst: 05:20
- L1 hf ramp burst: 09:00, 16:45
1.1194 - 14:52 (0.6pa 004)
- 18:54
(1.2pa 220)
- 22:35
(0.2pa 321)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.845 99.856 99.836 99.845 x 99.425 99.955 99.591 99.430   Ultisat
- Dwight: battery pack switch not on
- Backup modem working updated

Daily log
- All: 63-65,71
- Signal analysis on specific events requested by michael
3/14/2014 73 - calib values set to 0 - H4 BDA 20s spikes
- L2 ramp burst: 14:00
- L4 hf ramp burst: 06:00
- L1 hf ramp burst: 11:10
1.3743 - 02:42 (0.1pa 194)
- 06:41
(0.1pa 001)
- 17:03
(0.3pa 339)
- 18:16
(0.2pa 257)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.856 99.867 99.847 99.856 x 99.433 99.955 99.596 99.438   Signal analysis
- Finish for data requested by michael, send email with all attachments and explanations
- Send michael updated spreadsheet at end of day

Soh script
- Modify script to create larger year plots with gridlines

Daily log
- All: 66-70
3/15/2014 74 - calib values set to 0 - H4 BDA 20s spikes
- L4 hf ramp burst: 09:40
1.3921 - 01:18 (0.1pa 049) - elec noise: L2, L4
- wnrs noise: L1,L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.866 99.876 99.858 99.866 x 99.441 99.956 99.602 99.446  
3/16/2014 75 - calib values set to 0 - H4 BDA 20s spikes
- L2 hf ramp burst: 00:15
- L4 hf ramp burst: 11:00
1.2183 - elec noise: L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.874 99.883 99.867 99.874 x 99.448 99.956 99.607 99.453  
3/17/2014 76 - calib values set to 0 - H4 BDA 20s spikes
- L2 hf ramp burst: 01:50
1.9243 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- L2 BDF out of range
- H3 BDF out of range
  100.000 100.000 100.000 100.000 x 99.882 99.890 99.874 99.882 x 99.455 99.957 99.612 99.460   Daily log
- Re-process: 74, 75
- All: 72-74

Glen – radio response
- Create telemetry map
- Create radio details spreadsheet
- Ask dick for details on his radio
- Meet with glen, send spreadsheet and map

Ctbto access
- Create new login
- Test services
- Discuss with randy on phone
- Request additional services
3/18/2014 77 - calib values set to 0 - H4 BDA 20s spikes 1.7991 - 16:55 (0.3pa 310)
- 18:52
(0.1pa 076)
- 19:24
(0.5pa 327)
- 19:41
(0.1pa 106)
- 19:50
(1.6pa 322)
- elec noise: L2
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all except H1 clk diff too large   100.000 100.000 100.000 100.000 x 99.888 99.896 99.881 99.888 x 99.462 99.958 99.617 99.467   Ctbt sso access
- Test, find that requests are still obtaining approval
- Respond to Svetlana that password access has been restored

Soh script
- Research java heap space fixes

Server
- Failed again at 9:30
- Changed automount settings, restart computer, several attempts to fix
- Email say-ho, work with him on phone – might be caused by another user attempting to connect
- Take offline for several hours
- Bring back online temporarily with afp at end of day

Nviar data
- Respond to michael’s questions
- Contact david re: getting data

Wnrs
- Start work on prefab for 10’ square test

Daily log
- All: 076
3/19/2014 78 - calib values set to 0 - H4 BDA 20s spikes
- L2 ramp burst: 12:00, 15:40
- L4 hf ramp burst: 07:30
- H4 hf ramp burst: 10:30
- L1 hf ramp burst: 09:50, 16:40
1.2172 - 18:35 (0.2pa 259)
- 19:28
(0.2pa 340)
- 19:30
(1.0pa 204)
- elec noise: H4, L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.894 99.902 99.888 99.894 x 99.469 99.958 99.622 99.474   Nviar data
- Respond to david with IP I would like to ftp from (hal)
- Write scripts and start moving all old nviar data
- Complete move and inform michael

Daily log
- All: 077

Prep for tomorrow’s pinon trip
- Print notes
- Load equipment
- Fuel truck
3/20/2014 79 - Install wnrs squ test at H2 - calib values set to 0 - H4 BDA 20s spikes
- L2 ramp burst: 12:00, 15:10
- L4 hf ramp burst: 07:00
- L1 hf ramp burst: 08:40, 16:40
1.1537 - 18:19 (0.1pa 236)
- 18:54
(0.2pa 239)
- 22:47
(0.6pa 237)
- 23:10
(0.2pa 236)
- elec noise: L1, L2
- wnrs noise: H1, L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- visit related flags
  100.000 100.000 100.000 100.000 x 99.899 99.907 99.893 99.899 x 99.476 99.959 99.627 99.481  
3/21/2014 80 - calib values set to 0 - H4 BDA 20s spikes
- L4 hf ramp burst: 12:30
- L1 hf ramp burst: 13:30, 15:45
- H4 hf ramp burst: 13:35
1.402 - 22:09 (0.5pa 263)
- 22:48
(0.2pa 259)
- elec noise: L1
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.904 99.911 99.898 99.904 x 99.483 99.959 99.632 99.487   Soh script
- Stuck on java heap memory problem again – restart and reprocess yesterday

Field return
- Unload truck
- Write report

Daily log
- All: 78, 79

Nvar data
- Continue working on ftp .w file download
3/22/2014 81 - calib values set to 0 - H4 BDA 20s spikes 1.8817 - elec noise: L4
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.909 99.915 99.903 99.909 x 99.489 99.960 99.636 99.494  
3/23/2014 82 - calib values set to 0 - H4 BDA 20s spikes
- L4 hf ramp burst: 11:00
- L1 hf ramp burst: 13:40, 15:15
1.3752 - elec noise: L1, L4
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all except H2 clk diff too large   100.000 100.000 100.000 100.000 x 99.913 99.919 99.907 99.913 x 99.495 99.960 99.641 99.500  
3/24/2014 83 - calib values set to 0 - H4 BDA 20s spikes
- L2 hf ramp burst: 00:45
- L4 hf ramp burst: 14:00, 15:30
1.3105 - 22:29 (0.3pa 282)
- 23:09
(0.2pa 281)
- elec noise: L2, L4
- wnrs noise: L1,L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.916 99.922 99.911 99.916 x 99.501 99.961 99.645 99.506   Daily log
- All: 80-82

Nvar data
- Continue working on ftp .w downloads

Randy
- Discuss authentication key procedures
- Work on ctbt web portal requests

Michael status call
- Discuss authentication tests coming up
- Wants nviar data for meteor day

Prep for tomorrow’s trip
- Load equip
- Fuel truck
3/25/2014 84 - connected H2 SQU
- fixed H3, L2 VBA
(accidental restart of H3)
- troubleshoot L4 VBA
- NDC offline
- calib values set to 0
- H4 BDA 20s spikes
- L2 hf ramp burst: 01:30
- L4 hf ramp burst: 13:50, 15:50
- 19:46 (0.6pa 222)
- 20:22
(1.0pa 225)
- N/A - all clk diff too large
- 1450 frames from I57 to NDC missing
- site visit related flags
  99.988 99.993 100.000 66.366 x 99.919 99.925 99.915 98.574 x 99.507 99.961 99.649 99.111   H3
- Fix cabling to bring VBA online
- Calib measurement: 17:18 – 14.02
- 14.02V/1422cnts = .009859

L2
- Fix cabling to bring VBA online
- Calib measurement: 18:06 – 13.96
- 13.96V/1458cnts = .009575

L4
- Attempt to fix cabling to bring VBA online.
- After several tests concluded that Mass position #1 is not working
- Observed many chew marks on coax near vault entry
3/26/2014 85 - attempt to fix zero calib, 2 software stop/start all's - NDC large backfill
- calib values set to 0
- H4 BDA 20s spikes 2.8327 - 03:32 (0.8pa 236) - elec noise: L3
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all except H1 clk diff too large
- 1 frame from I57 to IDC delayed
- 2 frames from I57 to IDC duplicated
- 4717 frames from I57 to NDC delayed
- 736 frames from I57 to NDC backfilled
- 80 frames from I57 to UCSD missing
- 4 frames from profile2 to UCSD delayed
- 98 frames from profile2 to UCSD missing
  97.863 97.682 97.824 97.863 x 99.840 99.839 99.834 98.547 x 99.488 99.934 99.628 99.097   Trip return
- Unload truck
- Trip report
- Update daily log

Daily log
- All: 83 – 84
- PFO currently experiencing high winds and stormy conditions

Nvar
- Continue ftp downloads

Zero calibs
- All sites started sending zeros in the calib field on March 11th except L1
- On March 11th we reconfigured all digitizers for the extra VBA channel
- In July 2013 we discussed this problem with all when it happened to H4 digitizer using smart24config_1.29 – the solution was to upgrade to smart24config_1.32
- Ask randy if he wants me to submit OR and wait or submit PR and fix now
- 16:30 -> stop all on both systems
- resend configuration to all digitizers with 1.32
- update smartdbconfig on profile2 to expect L4 VBA from mass position 2 instead
- update L4 digitizer to send mass position 2 instead
- confirm all sensitivies as expected via telnet low-level ssl query’s
- confirm all calibs are as expected via telnet low-level ccl query’s
- umoya text files still show 0 calib values?
- Sent in report with update on reason for reboots
- Asked gary for help
- Update all on progress

Vba
- Calculate ncalibs from data collected yesterday
- Moved L4 VBA channel from MP #1 to MP #2 (see above)
- Updated soh script to input these new values
- They are wrong –> need to work on!
3/27/2014 86 - calib values set to 0 - H4 BDA 20s spikes 2.898 - elec noise: L2, L3
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- all BDF out of range
- 2745 frames from I57 to NDC delayed
- 634 frames from I57 to NDC backfilled
  100.000 100.000 100.000 100.000 x 99.846 99.845 99.840 98.601 x 99.494 99.935 99.632 99.107  
3/28/2014 87 - calib values set to 0 - H4 BDA 20s spikes
- L2 ramp burst: 06:35, 16:15
- L1 hf ramp burst: 17:30
1.6242 - 19:03 (0.6pa 192) - elec noise: L2
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.851 99.850 99.846 98.651 x 99.500 99.936 99.636 99.117  
3/29/2014 88 - Profile2 stopped receiving L2 data at 18:34
- calib values set to 0
- H4 BDA 20s spikes
- L4 hf ramp burst: 06:30
- L1 hf ramp burst: 09:50, 15:55
1.2285 - 22:45 (0.1pa 276)
- 23:21
(0.2pa 243)
- 23:27
(0.1pa 243)
- elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all except H3 clk diff too large
- L2 BDF, BDA missing data
  100.000 98.585 100.000 100.000 x 99.857 99.807 99.851 98.697 x 99.505 99.920 99.640 99.127  
3/30/2014 89 - Profile2 data stopped at 18:32
- calib values set to 0
- H4 BDA 20s spikes 3.0235 - 01:55 (0.1pa 245) - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all except H2 clk diff too large
- 982 frames from profile2 to UCSD missing
  100.000 72.418 100.000 100.000 x 99.861 98.894 99.856 98.741 x 99.511 99.611 99.644 99.137   Soh script
- Has been breaking all weekend, first because AFP server stopped working so switched over to NFS. Then started breaking because Profile2 not receiving L2
- Modified “run_previous” and backfilled missing days where could
- Turned of cronjob
3/31/2014 90 - profile2: stopall, clear queues, fix tables, startall - Profile2 restored at 13:43
- calib values set to 0
- H4 BDA 20s spikes
- L2 ramp burst: 12:00, 15:15
- L1 hf ramp burst: 16:30
2.339 - elec noise: L2
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all except H2 clk diff too large
- 22 frames from I57 to NDC delayed
- 2472 frames from profile2 to UCSD missing
  100.000 42.824 100.000 100.000 x 99.866 97.085 99.861 98.781 x 99.516 98.980 99.648 99.147   Profile2 outage:
- Stopped receiving L2 at:
- When logged on this morning the forwarder had stopped
- 13:44 - Cleared all forwarder files, repaired tables and restarted

Daily log
- Re-process: 88, 89
- Backfill missing smiar/mriar data
- Backfill fullyear graphs

Pinon Radios
- Inform glen that I can help with radio test, discuss scheduling

Pinon calibs
- Download H1, L1 config files and latest receiver log – send all to Gary

Michael status call
- Ndc outage
- Profile2 database corruption
- Pinon calibs
- Nvar data
- Authentication, return token
- Socal range complex?
- Inform ctbt of drilling
- Michael ask frank about drilling on Thursday morning during launch

Soh script
- Update to cat receiver files before parsing so that it includes all different receiver logs even when it’s restarted.
- Also had to modify to operate on log that is copied over to filel2a since matlab isn’t allowed to change umoya files directly

VBA calibs
- They are ~ 20x smaller than should be
- Looked and dbpick was applying some other calib to the “cts”, reset and got correct cts values to recalculate calips
- Updated soh script with new calibs
                    x x x x x x x x x x x x x x                                                                                                                                                  
4/1/2014 91 - L2 hf ramp burst: 07:25 2.7909 - elec noise: L2
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.522 98.992 99.652 99.156   Daily log
- Soh script failed due to java error, restart & reprocess yesterday

Timesheet
- Fill out spreadsheet based on daily log
- Email michael with summary and request for ok

VBA calibs
- Update fullyear_batch script to also apply new calibs and run for past 90 days

Ctbto portal
- Speak with randy on phone, would like me to test token
- Attempt to initialize token, does not work
- Sent message to ctbt support requesting assistance
- Discuss results with randy on phone

Soh script
- Fix fullyear so that it doesn’t use as much java memory
- Fix plots to fit data more accurately
- Fix receiver log file process so it better handles multiple receiver files and name changes

Zero calibs
- Ask jay if he has seen something similar

Nvar
- Wrote nviar_process script and batch script to process all of the data already downloaded
- Then added to nviar_process script to automatically ftp data from smu
- Then added nviar_process to soh_select_day so that entire process happens within the soh script at end of each day
- All files are currently up to date
4/2/2014 92 - H4 BDA 20s spikes
- L2 ramp burst: 12:30
- L2 hf ramp burst: 04:45
1.9792 - elec noise: L2
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.527 99.003 99.656 99.165   Soh script
- Stuck on end of day again, made changes to fullyear and re-ran
- Modify nviar_process to also cp .site into each day
- Write new function to reprocess I57 data over profile2 data on days where profile2 data is corrupted

Nvar
- Ask david for calib values
- Backfill .site for all ftp data
- Email michael and inform all data is ready

Monthly report
- File emails

Daily log
- All: 83 – 89
- Reprocess I57 data over profile2 data for 88-90
4/3/2014 93 - recorded Vandenberg launch - H4 BDA 20s spikes
- L2 ramp burst: 03:50, 17:50, 18:50
- H3 hf ramp burst: 15:30
1.5583 - 15:10 (0.5pa 287)
- 17:03
(0.4pa 329)
- 17:39
(0.7pa 265)
- 18:29
(1.6pa 271)
- 19:00
(1.2pa 261)
- 22:48
(1.0pa 244)
- elec noise: H3, L2
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.532 99.013 99.660 99.174   MR - 101674 Daily log
- All: 90-92

Vandenberg
- Monitor live webcast – successful launch at 17:46
- Analyze data – good recording of launch at 15:10
- Send michael pmcc and dbpick

Nviar
- Did not automatically process last night because files were not available early enough
- From the file times it looks like they post to ftp server at 03:05UT everyday

Soh script
- Modified soh_select_day to process the nviar data after 04:00 to make sure smu has it posted
- Modified process script to apply appropriate calibs to H2 SQU channel

MR
- Submit March report
4/4/2014 94 - H4 BDA 20s spikes
- L2 ramp burst: 06:30, 16:25
- L4 hf ramp burst: 03:30
- L1 hf ramp burst: 05:40
1.4009 - 18:45 (0.2pa 293)
- 19:16
(0.2pa 243)
- 19:32
(0.2pa 331)
- elec noise: H1, L1, L2
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.537 99.024 99.663 99.183   PR - 101469 Calibs
- Look through previous logs, write response to ctbt with history
4/5/2014 95 - H4 BDA 20s spikes
- L2 ramp burst: 07:05, 13:40
- L1 hf ramp burst: 18:20
2.4083 - 18:52 (0.5pa 257) - elec noise: L2
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.542 99.034 99.667 99.192  
4/6/2014 96 - H4 BDA 20s spikes
- L2 ramp burst: 11:50, 14:50
- L1 hf ramp burst: 06:50
1.5188 - 03:28 (0.3pa 243) - elec noise: L1, L2
- wnrs noise: H2, L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.546 99.044 99.670 99.200  
4/7/2014 97 - H4 BDA 20s spikes
- L2 ramp burst: 19:00
- L4 hf ramp burst: 11:00
1.4478 - 00:53 (0.1pa 243)
- 17:19 (0.1pa 025)
- 19:53
(0.1pa 342)
- elec noise:
- wnrs noise: H2, L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 2 frames from I57 to IDC delayed
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.551 99.054 99.674 99.208  
4/8/2014 98 - H4 BDA 20s spikes
- L2 hf ramp burst: 01:50
- H1 hf ramp burst: 23:10
1.2893 - 16:51 (0.3pa 207)
- 21:30
(0.2pa 252)
- elec noise: L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 9 frames from I57 to IDC delayed
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.556 99.064 99.677 99.217  
4/9/2014 99 - H4 BDA 20s spikes
- H2 hf ramp burst: 01:40
- H4 hf ramp burst: 00:10, 02:50
- H1 hf ramp burst: 03:00
- L2 hf ramp burst: 03:10
1.2726 - elec noise: H2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all except L3 clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.560 99.073 99.680 99.224  
4/10/2014 100 - Glen installed 6 Vernon radios - H4 BDA 20s spikes
- H1 hf ramp burst: 04:10
- H2 hf ramp burst: 03:35
- H4 hf ramp burst: 04:00
- L2 hf ramp burst: 03:50
1.5319 - 23:13 (0.2pa 277) - elec noise: L1-L3
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all except H3 clk diff too large
- 3 frames from Profile2 to UCSD delayed
- 2 frames from Profile2 to UCSD duplicated
- 33 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.565 99.082 99.684 99.232  
4/11/2014 101 - Glen installed 4 Vernon radios - H4 BDA 20s spikes
- H4 hf ramp burst: 04:10
- H1 hf ramp burst: 04:50
- H2 hf ramp burst: 04:40
1.8532 - elec noise: L1-L3
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 33 frames showing H1 GPS unlocked
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.569 99.091 99.687 99.240  
4/12/2014 102 - H4 BDA 20s spikes
- BWS error spike: 15:10
- H4 hf ramp burst: 04:40
2.1625 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 2 frames from Profile2 to UCSD delayed
- 1 frame from Profile2 to UCSD duplicted
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.573 99.100 99.690 99.247  
4/13/2014 103 - H4 BDA 20s spikes
- L1 hf ramp burst: 14:20
- H4 hf ramp burst: 22:00
2.4884 - elec noise: L1
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.577 99.109 99.693 99.255  
4/14/2014 104 - H4 BDA 20s spikes
- H4 hf ramp burst: 03:05
- H1 hf ramp burst: 04:10
1.5737 - 20:55 (0.6pa 274)
- 21:29
(1.5pa 274)
- 22:03
(1.0pa 273)
- 22:37
(0.3pa 276)
- 23:08
(0.3pa 271)
- 23:29
(0.2pa 273)
- elec noise:
- wnrs noise: L1-L2
- 8hz bump: L1-L4
- all clk diff too large
- 55 frames from I57 to NDC delayed
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.581 99.118 99.696 99.262  
4/15/2014 105 - H4 BDA 20s spikes
- H1 ramp burst: 03:45
- H4 hf ramp burst: 01:30, 23:00
- H2 hf ramp burst: 10:10
- L1 hf ramp burst: 11:30, 15:40
1.2537 - elec noise: H4, L1
- wnrs noise: H2, L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.585 99.126 99.699 99.269  
4/16/2014 106 - Generated authentication keys for all - H4 BDA 20s spikes 1.7687 - elec noise:
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 37 frames from I57 to IDC missing
- 37 frames from I57 to NDC missing
- 37 frames from I57 to UCSD missing
  99.144 100.000 99.144 99.144 x 99.946 100.000 99.947 99.946 x 99.581 99.134 99.693 99.268   Timesheet
- Submit

Pki portal
- Successfully initialize token
- Successfully login to access.ctbto.org
- Unable to login to pki portal
- Email randy: get correct login name
- Successfully login to pki portal
- Respond to ctbto re: successful token initialization
- Email Jay: did he stop software before key generation

Generate keys:
- Stopall - 18:40:00
- Generate all keys
- Startall
- Edit keys and upload to pki portal
- Inform all

Vernon – pinon radios
- Monitor packet loss – L1 losing 1.25 packet/min
- Discuss with michael
- Email glen and all

Research linux computer for ssi
4/17/2014 107 - H4 BDA 20s spikes
- H2 hf ramp burst: 02:10
- H1 hf ramp burst: 05:20, 19:35
- H4 hf ramp burst: 04:40, 22:40
1.305 - elec noise: H4
- wnrs noise: H1, L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.950 100.000 99.950 99.950 x 99.585 99.142 99.696 99.274   Ssi computer
- Will run CentOS version x.x?
- Want Ipmi – monitor system voltage, fans, etc. : control computer outside OS
- Want hot swapable
- Rack mount?
- Redundant power supply, digital switching?
- Ask jay which version and if card reader is working on it
- Good interface with new UPS
- Sent request for quote to local vendor
- Long email to michael explaining all

Authentication
- Sent key to isla
4/18/2014 108 - H4 BDA 20s spikes
- H1 ramp burst: 03:30
- H4 hf ramp burst: 03:05
- H1 hf ramp burst: 20:40
1.3372 - elec noise: H1, L3
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.952 100.000 99.952 99.952 x 99.589 99.150 99.699 99.281   Michael status call:
- He will send curt email with battery, ups and ssi computer costs
4/19/2014 109 - H4 BDA 20s spikes
- BWS error spike: 13:26
- L4 hf ramp burst: 12:20, 16:00
- L2 hf ramp burst: 14:00
- L1 hf ramp burst: 15:00
1.0622 - 00:09 (0.3pa 236) - elec noise: L1, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.955 100.000 99.955 99.955 x 99.593 99.158 99.702 99.288  
4/20/2014 110 - H4 BDA 20s spikes
- H1 hf ramp burst: 22:40
1.2868 - elec noise: H1
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.957 100.000 99.957 99.957 x 99.596 99.166 99.704 99.294  
4/21/2014 111 - H4 BDA 20s spikes
- H1 hf ramp burst: 03:20
- H4 hf ramp burst: 02:35, 21:40
- L2 hf ramp burst: 03:40
1.3389 - 06:36 (0.3pa 001)
- 11:06
(0.1pa 003)
- elec noise: H1, H4, L4
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 33 frames from I57 to NDC delayed
  100.000 100.000 100.000 100.000 x 99.959 100.000 99.959 99.959 x 99.600 99.173 99.707 99.301   Daily log
- All 93 – 94

Michael status call

Signal analysis:
- research point mugu
- Vandenberg launch analysis
4/22/2014 112 - H4 BDA 20s spikes 2.1207 - elec noise: L3
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.961 100.000 99.961 99.961 x 99.604 99.181 99.710 99.307   Daily log
- All: 95 – 102

Ssi
- Respond to ryan’s questions

Michael status meeting
– discuss all
- Give michael april timesheet estimate
4/23/2014 113 - start authentication keys - H4 BDA 20s spikes
- H1 hf ramp burst: 21:20
2.0703 - 18:37 (0.1pa 110)
- 19:02
(0.1pa 108)
- 22:20
(0.1pa 109)
- elec noise: H1
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 28 frames from I57 to IDC missing
- 28 frames from I57 to NDC missing
- 28 frames from I57 to UCSD missing
- H1 BDF out of range
- H4 BDF out of range
  99.352 100.000 99.352 99.352 x 99.935 100.000 99.935 99.935 x 99.601 99.188 99.707 99.307   Milton call
- Wanted to know if I received any training for the key generation

Authentication key
- Ask randy what format I use for responding after starting keys
- Discuss with david in Hawaii – they are at same spot
- Start keys:
- 21:30 – stop all
- Start all keys, confirm forwarder and server config files have signing to idc turned on
- 21:40 – start all
- Respond to pts with start times
- Send message to group explaining informal response

Ssi computer
- Spend most of day researching fanless ssd computers, narrow down to 4 different manufacturer/models
- Respond to ryan’s questions for supermicro computer quote
- Send update to randy with prices I’ve found so far
4/24/2014 114 - H4 BDA 20s spikes
- H1 hf ramp burst: 03:20, 21:50
- H4 hf ramp burst: 14:00
- L1 hf ramp burst: 14:10
1.2571 - 23:30 (0.2pa 293) - elec noise: L1
- wnrs noise H1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.937 100.000 99.937 99.937 x 99.605 99.195 99.709 99.313   Daily log
- All: 103-104
- Dbpick research – quick commands

Ssi
- Receive quote from ryan
- Continue research on pro/cons of each system
- Email michael with prices/quotes and long description of choice
4/25/2014 115 - H4 BDA 20s spikes
- H1 hf ramp burst: 03:45
2.1674 - 00:25 (0.2pa 195) - elec noise: H1
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.940 100.000 99.940 99.940 x 99.608 99.202 99.712 99.319   Ctbt product
- Received 112 emails on last night’s stats
- Sent message to support informing them

Randy:
- They accepted my email and my keys are being authenticated
- Auth_id field is certificate serial number – open ssl

Authentication
- Pts sent another request for us to populate the auth_id field
- Researched
- Discussed with gary
- Email group asking for comments
4/26/2014 116 - H4 BDA 20s spikes 2.9507 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- H1-H4 BDF out of range
  100.000 100.000 100.000 100.000 x 99.942 100.000 99.942 99.942 x 99.612 99.209 99.714 99.325  
4/27/2014 117 - H4 BDA 20s spikes
- L2 ramp burst: 15:25
- L1 hf ramp burst: 17:00
2.1525 - elec noise: L1
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.944 100.000 99.944 99.944 x 99.615 99.216 99.717 99.331  
4/28/2014 118 - H4 BDA 20s spikes 1.6589 - elec noise:
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.946 100.000 99.946 99.946 x 99.618 99.222 99.719 99.337   Randy/gary call
- Clearing up confusion – it’s been called cert_id and auth_id but geotech refers to it as key_id and key_id is the only field they can change
- The way it “should” work is we generate a key, the pts sends back a certificate (which contains the key_id in it. We put that certificate on the workstation and the start_key process reads that certificate, pulls out the key_id and uploads it to the digitizer during the start_key process.
- We asked gary if we could have a quick work around and just upload the new key_id to the digitizer manually without generating a new key.

Daily log
- All: 105 – 113
4/29/2014 119 - H4 BDA 20s spikes
- H1 hf ramp burst: 00:20, 01:20, 02:20
2.0008 - 22:32 (0.3pa 019)
- 22:54
(1.2pa 229)
- elec noise: H1, L2
- wnrs noise: H2, L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 9 frames from Profile2 to UCSD delayed
- 5 frames from Profile2 to UCSD duplicated
  100.000 100.000 100.000 100.000 x 99.948 100.000 99.948 99.948 x 99.621 99.229 99.721 99.342   Daily log
- All: 114-118

Soh script
- Updated fullyear to overlap VBA plots
- Update naming scheme for all files of program
- Start work on new vba data in soh section

Proposal
- Gave michael precise quotes on UPS, computer and batteries
4/30/2014 120 - H4 BDA 20s spikes
- BWS error spike: 01:04
- H1 hf ramp burst: 02:10
2.4238 - elec noise:
- wnrs noise: H1-H2, L1-L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.950 100.000 99.950 99.950 x 99.625 99.235 99.724 99.348   Daily log
- All: 119

Gary call
- Calib sensitivity issue
- Send certificate that pts gave me
- Low level command for getting key “ark” will return the key_id -> Ark (0 or 1) effective or pending key

Authentication
- Update randy on all learned, ask if pts can give keys first

Wnrs
- Purchase supplies for next phase of square filter at home depot
- Submit receipt

Tantor
- Update main function by changing button callbacks and creating new smaller independent functions in another folder rather than having all callbacks in main function

Inventory
- Respond to maria/michael re: ucsd inventory clairification request
                    x x x x x x x x x x x x x x          
5/1/2014 121 - H4 BDA 20s spikes
- H1 hf ramp burst: 21:40
- H4 hf ramp burst: 23:40
2.4628 - 01:48 (0.2pa 238)
- 02:12
(0.1pa 022)
- 05:26
(0.1pa 276)
- elec noise: H4
- wnrs noise: H2, L1-L4
- 8hz bump: L1-L4
- all clk diff too large
- 2354 frames from I57 to NDC missing
  100.000 100.000 100.000 45.370 x 100.000 100.000 100.000 45.370 x 99.628 99.242 99.726 98.902   Daily log
- All: 120

Timesheet:
- fill out, copy, submit

Randy call
- Discuss key_id possibilities
- Discuss 111 email from pts

Authentication
- Give randy/gary the key_id’s I used
- Contact spyrus – leave msg requesting pricing/availability on new ssi authenticator

Calibs
- Give everybody history
- Confirm that I left sensitivity field blank
- Ask jay to provide his sensitivity values

Tantor
- Complete callback modification to main script
- Update limit table with new vba info

Ssi
- Asked randy if we will need to purchase the spyrus authenticator jay mentioned
- Spoke with spyrus:
• Cost is $800 ($1360 with tax, ship, o.h.)
• Lead time 3 weeks
• Does not work with 64-bit
• The mercury system is 64-bit
- Informed randy/michael of all
- Gave randy our card reader model number
- Asked jay about his system and learned that we need to go with 32-bit for this and centos reasons
- Ask ryan if he can revise quote for a 32-bit system
5/2/2014 122 - H4 BDA 20s spikes
- H2 hf ramp burst: 00:20, 02:30, 20:20
- H1 hf ramp burst: 04:10, 18:50
- H4 hf ramp burst: 03:10, 20:30
- L2 hf ramp burst: 03:15
1.3938 - 17:56 (0.2pa 236)
- 20:24 (0.2pa 229)
- elec noise: H4, L2, L4
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 5113 frames from I57 to NDC delayed
- 2354 frames from I57 to NDC backfilled
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 72.685 x 99.631 99.248 99.728 98.911   Authentication
- Several discussions with randy re: a system for us providing unique key_id’s
- Still not sure if spyrus authenticator will be necessary or not

Proposal
- Provide several documents to michael
- Research available figures

Daily log
- All: 121

Batteries
- Place order for 8 batteries through mouser
5/3/2014 123 - H4 BDA 20s spikes
- H1 hf ramp burst: 06:40
- H2 hf ramp burst: 05:20, 17:20
- H4 hf ramp burst: 05:35, 18:15
- L2 hf ramp burst: 05:00
1.62 - elec noise: H4
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all except H3 clk diff too large   100.000 100.000 100.000 x 100.000 100.000 100.000 81.790 x 99.634 99.254 99.728 98.919  
5/4/2014 124 - H4 BDA 20s spikes
- H4 hf ramp burst: 07:05, 17:45
- H1 hf ramp burst: 11:00
1.9442 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 x 100.000 100.000 100.000 86.343 x 99.637 99.260 99.728 98.928  
5/5/2014 125 - H4 BDA 20s spikes 2.1953 - elec noise: H1
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all except H2 clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 89.074 x 99.640 99.266 99.730 98.937   PR - 90782
MR - 102384
Tantor
- Wrote new “reprocess” script to fill in lost days
- Convert fullyear from script to function

Calibs
- Research and write long response to Svetlana report detailing what the calib values should be and what actions are needed

Daily log
- Reprocessed 122-124

Submit MR
5/6/2014 126 - H4 BDA 20s spikes 2.1058 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 4 frames from I57 to IDC delayed
- 2 frames from I57 to IDC duplicated
- 1 frame from I57 to NDC duplicated
- 1 frame from I57 to UCSD duplicated
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 90.895 x 99.643 99.272 99.733 98.945   Wnrs
- Build pipes
- Load truck
- Fuel truck

Daily log:
- All: 122
5/7/2014 127 - restart profile2 software at 15:35 - Profile2 experienced server port bind error with L3 port - H4 BDA 20s spikes
- L2 ramp burst: 07:30
- L2 hf ramp burst: 15:20
- N/A - all clk diff too large
- 3 frames from profile2 to UCSD delayed
- 10 frames from profile2 to UCSD missing
  100.000 99.222 100.000 100.000 x 100.000 99.889 100.000 92.196 x 99.645 99.271 99.735 98.953   Ssi computer
- Email all with details to work out

Randy call:
- Calibs:
- Do I proceed or wait for response -
• BDA/BDF still won’t work because of mp issue
- Ssi:
• 32bit vs 64bit -
• Fanless vs server -
• Spare for server -
• Authenticators now –
- Reports fixed
- Order placed for L3/L4 batteries
- PFO calibration
- Key_id

L3 offline
- Happened at 13:45
- Could be due to batteries, happened just before sunrise, voltage was at 11.9 (although H1 seems ok at 11.4)
- I57 computer does not show issue with L3
- Smart24config can log on to L3 and does not show issue
- Profile2 showing port bind error on smartserver
- 15:35 – stopall, clear queues, repair tables, startall -> Working

Wnrs
- Purchase more supplies at home depot
- Finish building prep pipes and manifolds
- Load up for trip tomorrow
5/8/2014 128 - upgrade 32-square to a 64-square
- repair hpwren switch power strip
- hpwren switch damaged - H4 BDA 20s spikes
- L2 ramp burst: 12:45, 14:15
- L1 hf ramp burst: 17:45
- H1 hf ramp burst: 21:30
1.7073 - 19:28 (0.1pa 359)
- 19:45
(0.1pa 002)
- 20:06
(0.2pa 048)
- 20:27
(0.2pa 108)
- elec noise: L2
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all except H1 clk diff too large
- 324 frames from I57 to NDC delayed
- 323 frames from I57 to UCSD delayd
- 328 frames from profile2 to UCSD delayed
- 2 frames from profile2 to UCSD duplicated
- site visit related flags
  100.000 100.000 100.000 100.000 x 100.000 99.903 100.000 93.171 x 99.648 99.277 99.737 98.962   All
- Visual inspection of exterior at all sites
5/9/2014 129 - H4 BDA 20s spikes
- H1 hf ramp burst: 03:50
- L1 hf ramp burst: 11:00, 15:10
1.6244 - 22:06 (0.4pa 203) - elec noise: L1
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 29 frames from I57 to NDC delayed
- 29 frames from I57 to UCSD delayed
- 49 frames from profile2 to UCSD delayed
- 4 frames from profile2 to UCSD duplicated
  100.000 99.977 100.000 100.000 x 100.000 99.911 100.000 93.930 x 99.651 99.282 99.739 98.970   Field return
- Unload truck
- Trip report

Daily log
- All: 127-128

Signal analysis
- Work on matlab filters, spec comparison scripts, etc…
5/10/2014 130 - hpwren switch offline intermittently - H4 BDA 20s spikes
- H4 hf ramp burst: 00:15, 01:50
- H1 hf ramp burst: 04:40
2.0254 - elec noise:
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 2364 frames from I57 to NDC delayed
- 2364 frames from I57 to UCSD delayed
- 2373 frames from profile2 to UCSD delayed
- 5 frames from proflie2 to UCSD duplicated
  100.000 100.000 100.000 100.000 x 100.000 99.920 100.000 94.537 x 99.654 99.288 99.741 98.978  
5/11/2014 131 - hpwren switch offline intermittently - H4 BDA 20s spikes
- L1 hf ramp burst: 12:10
2.4047 - elec noise:
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- BDF flags at all sites
  100.000 100.000 100.000 100.000 x 100.000 99.927 100.000 95.034 x 99.656 99.293 99.743 98.985  
5/12/2014 132 - hpwren switch offline intermittently - H4 BDA 20s spikes
- H1 ramp burst: 04:00
- L1 hf ramp burst: 11:50
- H4 hf ramp burst: 23:05
1.7395 - elec noise: L1, H4
- wnrs noise: H1, L1, L2
- 8hz bump: L1-L2
- all clk diff too large
- 2960 frames from I57 to NDC delayed
- 2956 frames from I57 to UCSD delayed
- 1315 frames from I57 to NDC backfilled
- 1319 frames from I57 to UCSD backfilled
- 2967 frames from profile2 to UCSD delayed
- 1321 frames from profile2 to UCSD backfilled
  100.000 100.000 100.000 100.000 x 100.000 99.933 100.000 95.448 x 99.659 99.299 99.745 98.993   PR - 90782
CCN - 102643
OR - 102644
Daily log
- Reprocess: 130 – 131 (due to hpwren switch outages)

Randy call
- Discuss calibs

Michael status call
- Weekend outages

Tantor
- Update to properly rename log files when we have receiver outages

Calibs
- Respond to Svetlana with new summary
- Submit ccn for BDA calib change
- Submit OR for BWD calib change
5/13/2014 133 - H4 BDA 20s spikes
- H1 ramp burst: 05:20
- H4 hf ramp burst: 03:30, 22:40
- L1 hf ramp burst: 11:10
1.8111 - elec noise: H1, H4
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 99.938 100.000 95.798 x 99.661 99.304 99.747 99.001   Daily log
- All: 132

Spec_pwr
- Write script that makes a comparison plot of h2 bdf vs squ
- Script also creates a matrix of the signals with psd ratios for each signal I analyze
5/14/2014 134 - H4 BDA 20s spikes
- H4 BDA offset after reconfigure
- H1 ramp burst: 05:30
- H2 hf ramp burst: 00:50
- H4 hf ramp burst: 04:00, 20:40
- L1 hf ramp burst: 12:10, 14:35
1.7167 - 06:00 (0.1pa 261)
- 20:17 (0.2pa 045)
- elec noise: H1, H2, H4, L1
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   99.799 99.758 99.560 99.799 x 99.986 99.925 99.963 96.084 x 99.662 99.307 99.746 99.007   PR - 101469 Calibs
- 14:37 – use ccl command to adjust L1 calib value
- 14:45 – 15:00 – use ccl command to adjust all other calib values
- Respond to PR and explain all

Daily log
- All: 133

Snr
- Continue work filling matrix
- Start work on normalization plot
5/15/2014 135 - H4 BDA 20s spikes
- H1 ramp burst: 06:25
- L2 hf ramp burst: 03:30
- H2 hf ramp burst: 04:30
- H4 hf ramp burst: 05:05
- L3 hf ramp burst: 05:05
1.4946 - 02:41 (0.1pa 354) - elec noise: H4
- wnrs noise: H2, L1, L2
- 8hz bump: L1-L4
- all except L3 clk diff too large
- 51 frames from I57 to IDC delayed
  100.000 100.000 100.000 100.000 x 99.987 99.930 99.966 96.345 x 99.665 99.313 99.747 99.014   Randy call
- Discuss the smart calib issue

Calibs
- Bring gary up to date

Michael status call
- he would like temperature sensors in all enclosures
- offer to put temp indicator labels in the vaults
5/16/2014 136 - H2 charge controller strangeness
- H4 BDA 20s spikes
- H1 ramp burst: 08:25
- H2 hf ramp burst: 06:35
- H4 hf ramp burst: 07:00, 16:40
- L2 hf ramp burst: 05:15, 09:40
- L3 hf ramp burst: 07:00
1.5278 - 03:12 (0.1pa 012) - elec noise: H4, L2, L3
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.987 99.935 99.969 96.573 x 99.667 99.318 99.749 99.021   Token:
- respond to Svetlana informing her the token is working

Daily log
- All: 134-135
- Reprocess: 135

Snr script
- Add signals
- Update with normalization routine
5/17/2014 137 H4 BDA 20s spikes 2.5402 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.988 99.939 99.971 96.775 x 99.670 99.323 99.751 99.028  
5/18/2014 138 - H4 BDA 20s spikes 2.5596 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all except L3 clk diff too large   100.000 100.000 100.000 100.000 x 99.989 99.942 99.973 96.954 x 99.672 99.327 99.753 99.035  
5/19/2014 139 - H4 BDA 20s spikes 2.5322 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all except H4 clk diff too large
- 2 frames from I57 to IDC delayed
- I57H4 BDF flags due to high winds
  100.000 100.000 100.000 100.000 x 99.989 99.945 99.974 97.114 x 99.675 99.332 99.755 99.042   Daily log
- Reprocess: 138

Batteries
- ask Helena why shipping cost has changed
5/20/2014 140 - updated BWD calib value - H4 BDA 20s spikes 2.1932 - 00:04 (0.8pa 111)
- 15:58 (0.2pa 111)
- elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   99.970 99.935 100.000 99.970 x 99.988 99.945 99.976 97.257 x 99.677 99.337 99.757 99.049   Daily log
- All: 139

Snr script
- Continue matlab work on xcorr to find and compare signals

BWD calib
- Following the OR times I updated the calib value at 17:00
5/21/2014 141 - install temperature tags, inspect vaults - H4 BDA 20s spikes
- L2 hf ramp burst: 09:35
2.016 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- site visit related flags
  100.000 100.000 100.000 100.000 x 99.989 99.947 99.977 97.388 x 99.679 99.341 99.758 99.056   L2
- Temperature stickers on batteries and digitizer

H1
- Temperature stickers on batteries and digitizer
- Confirmed batteries are no longer in dual chg cntrlr configuration, the wiring was left in place but disconnected, the batteries are in the normal config like all other sites
- 12:05 – 13.33v
- Significant deviation when calculated calib
- For the lid hasps the vault wall thickness is 3.75”
- The post to hold the lid would need to stick out of ground approximate 3’
- The seal should be centered on the vault lid as follows: 39.5” x 63.5”

H3
- Placed stickers on charger controller, digitizer, radio, dc converter
5/22/2014 142 - H4 BDA 20s spikes
- H1 hf ramp burst: 04:30
- L1 hf ramp burst: 09:50, 15:40
1.3708 - 16:02 (0.1pa 108)
- 17:28
(0.1pa 104)
- elec noise: H1, L1
- wnrs noise: L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.989 99.950 99.978 97.506 x 99.681 99.346 99.760 99.062   Michael signal analysis
- Provide info on all arrays for aerosmith study

Batteries
- Respond to mouser twice regarding shipping

Calibs
- Respond to david regarding work completed

Daily log
- All: 140-141

Trip return
- Write report – confirm results

I57H1 VBA calib
- Found from yesterday’s measurement that the H1 calib has been incorrect.
- Wrote script and ran to apply new calib to all data for days 91-current and to update the fullyear plots
5/23/2014 143 - bad frame causing fwrdr issues - H4 BDA 20s spikes
- L2 ramp burst: 10:20, 14:35
- L1 hf ramp burst: 05:50, 16:00
- L4 hf ramp burst: 10:50, 15:30
1.1715 - 19:38 (0.1pa 284) - elec noise: L2, L3
- wnrs noise: H1, L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 1 frame from I57 to IDC missing
- 5 frames from I57 to NDC delayed
- 1 frame from I57 to NDC duplicated
- 6 frames from I57 to UCSD delayed
- 1 frame from I57 to UCSD missing
- 6 frames from profile2 to UCSD delayed
- 1 frame from profile2 to UCSD missing
  99.954 99.977 99.977 99.977 x 99.988 99.951 99.978 97.614 x 99.683 99.350 99.762 99.069   Daily log
- 142: all

Ndc in a box
- Ask randy if he has heard back from pts on updated version

Wnrs
- Use autocad to design new plans for next square design: 64 inlet, 35’ square

Snr
- Write new script to better automate and include more background noise
5/24/2014 144 - H4 BDA 20s spikes
- H4 hf ramp burst: 01:30
- H1 hf rampburst: 05:00, 17:20
1.4712 - elec noise: H4
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all except L3 clk diff too large   100.000 100.000 100.000 100.000 x 99.988 99.953 99.979 97.713 x 99.685 99.355 99.763 99.075  
5/25/2014 145 - bad frame causing fwrdr issues - H4 BDA 20s spikes
- H1 ramp burst: 07:30
- H2 hf ramp burst: 05:40, 17:30
- H4 hf ramp burst: 06:20, 18:00
- H1 hf ramp burst: 16:40
1.4848 - elec noise: H2, H4, L3
- wnrs noise: H1, L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 1 frame from I57 to IDC delayed
- 2 frames from I57 to IDC duplicated
- 13 frames from I57 to IDC missing
- 130 frames from I57 to NDC delayed
- 7 frames from I57 to NDC duplicated
- 7 frames from I57 to NDC missing
- 115 frames from I57 to UCSD delayed
- 8 frames from I57 to UCSD duplicated
- 7 frames from I57 to UCSD missing
- 140 frames from profile2 to UCSD delayed
- 7 frames from profile2 to UCSD missing
  99.468 99.838 99.653 99.375 x 99.968 99.948 99.965 97.780 x 99.684 99.358 99.762 99.077  
5/26/2014 146 - H4 BDA 20s spikes
- L2 hf ramp burst: 07:10, 11:25
- H1 hf ramp burst: 10:50, 15:30
- H2 hf ramp burst: 10:20, 15:20
- H4 hf ramp burst: 15:00, 23:30
1.7682 - elec noise: H1, H4, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.969 99.950 99.966 97.865 x 99.686 99.363 99.764 99.084  
5/27/2014 147 - clear queue, repair tables, restart - H4 BDA 20s spikes
- H1 ramp burst: 12:20
- H4 hf ramp burst: 04:00, 04:45, 22:10
- L2 hf ramp burst: 09:40
- H1 hf ramp burst: 15:00
1.8732 - 17:00 (0.6pa 109) - elec noise: H1, H4, L2, L3
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 8 frames from I57 to IDC missing
- 8 frames from I57 to NDC missing
- 8 frames from I57 to UCSD missing
- 14 frames from profile2 to UCSD missing
  99.815 99.676 99.815 99.815 x 99.963 99.940 99.960 97.937 x 99.687 99.365 99.764 99.089   Daily log
- All:
- Reprocess: 145 – 146

I57 outage
- Tried several times over the weekend to restart tantor
- Monitored fwrdr over weekend and saw recorded frame gaps but sending seemed to be working
- On 5/23 at 17:55 the forwarder complained about an “errorred” data frame
- After that there are many backfill events for the day
- 5/24 seems ok but on 5/25 the errors return with several complains about socket write errors and many more backfill events
- On 5/26 the system is ok again with no reported errors however it is reporting several gaps.
- The strange part is it’s happening on both the primary I57 computer and the profile2 computer as well?
- Profile2 is showing the same gap records as I57.
- 15:52 – profile2 – clear queues, repair tables, restart
• Everything appears ok
- 16:00 – I57 – clear queues, repair tables, restart
• Working but still showing gap list, will let run for next couple of days to see if any more problems come back otherwise will have to perform complete database cleanup
- In order for both systems to have been affected we assume a digitizer sent a bad frame to both profiles:
• Nothing in either server log around 5/23 17:55 that shows a bad frame?

Michael status call
- Leave message

SNR
- Continue working on script
5/28/2014 148 - H4 BDA 20s spikes
- H4 hf ramp burst: 05:10
- L2 hf ramp burst: 11:30, 15:05
1.6461 - 15:36 (0.1pa 107)
- 17:15
(0.1pa 328)
- elec noise: H2, L3
- wnrs noise: H1, L1, L2
- 8hz bump: L1-L4
- all except L3 clk diff too large
- I57H2 BDF flag
  100.000 100.000 100.000 100.000 x 99.964 99.942 99.962 98.011 x 99.689 99.369 99.766 99.095   Daily log
- All: 147

SNR
- Continue working on script
5/29/2014 149 - H4 BDA 20s spikes
- H4 hf ramp burst: 22:20
2.263 - 16:04 (0.1pa 109) - elec noise:
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all except H2 clk diff too large
- 24 frames from I57 to IDC delayed
  100.000 100.000 99.977 100.000 x 99.966 99.944 99.962 98.080 x 99.691 99.373 99.768 99.101   Daily log
- All: 148

SNR
- Continue working on script
5/30/2014 150 - H4 BDA 20s spikes
- H4 hf ramp burst: 04:45, 22:15
- L2 hf ramp burst: 07:45
1.7827 - elec noise: L2, L3
- wnrs noise: H1, L1, L2
- 8hz bump: L1-L4
- all except H1 clk diff too large   100.000 100.000 100.000 100.000 x 99.967 99.946 99.964 98.144 x 99.693 99.377 99.769 99.107   Daily log
- All: 149

Snr
- Continue working on script
- Start inputing data

Receive batteries

Attempt ndc n a box dnld, inform randy it’s not working
5/31/2014 151 - H4 BDA 20s spikes
- L1 single spike: 08:30
- H4 hf ramp burst: 02:30, 05:25
- L2 hf ramp burst: 08:30
1.8511 - elec noise: L3
- wnrs noise: H1, L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.968 99.948 99.965 98.203 x 99.695 99.382 99.771 99.113  
                B   x x x x x x x x x x x x x x          
6/1/2014 152 - H4 BDA 20s spikes
- H4 hf ramp burst: 22:10
2.0715 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all except L2 clk diff too large
- L2 BDF flag
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.697 99.386 99.772 99.119  
6/2/2014 153 - H4 BDA 20s spikes
- L2 hf ramp burst: 07:10
2.1685 - elec noise:
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.699 99.390 99.774 99.124   Randy:
- Has new 64 bit ssi version

Daily log
- Reprocess: 150, 152

Batteries
- Unpack all, check all voltages
- Start charging

Timesheet
6/3/2014 154 - H4 BDA 20s spikes
- H4 hf ramp burst: 03:40
1.9924 - elec noise:
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all except H2 clk diff too large
- 1 frame from profile2 to ucsd duplicated
- L2, H1, H2 BDF flags due to high winds
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.701 99.394 99.775 99.130  
6/4/2014 155 - H4 BDA 20s spikes
- H4 hf ramp burst: 04:25, 12:30, 14:40, 22:45
- L2 hf ramp burst: 05:40
2.0331 - 14:34 (0.2pa 109)
- 17:32
(0.2pa 108)
- elec noise:
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all except H1 clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.703 99.398 99.777 99.136   MR - 103140 Express card
- Explain extra shipping to lenore and provide new receipt

Ssi
- Randy:
• Do I need to respond to ccr
• Am I correct that we don’t plan to use for next year
• Am I correct that we don’t purchase until funds from 1st bridge received

Randy call
- Engineer meeting march 15’ – training for ssi and secure portal
- Trillium is ok, carl contacted him
- Needs info on wnrs so they can build one

ccr
- Inform michael that no action is needed on our part
- Ask randy about email list

Daily log
- All: 153, 154

MR
- File emails
- Submit report

Snr
- Continue filling database
6/5/2014 156 - umoya cdrecv software restarts - H4 BDA 20s spikes
- H4 hf ramp burst: 04:30, 13:50, 21:15
- L2 hf ramp burst: 05:15
1.9849 - 14:44 (0.1pa 107)
- 21:58
(0.1pa 105)
- elec noise: H4
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all except H1 clk diff too large
- 1 frame delayed from profile2 to ucsd
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.705 99.401 99.778 99.141   Authentication
- Send randy the auth_key_id’s used

Batteries
- Continue charging, testing, inventory update

Umoya
- Attempt to fix “cdrecv shutdown” that occurs everytime I open the running window by changing the debug level from 9 to 1.
- Also noted the version: 2.1.34

Daily log
- All: 155
6/6/2014 157 - H4 BDA 20s spikes
- H4 hf ramp burst: 20:20
2.497 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- I57L1 BDF flag
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.707 99.405 99.780 99.147   Daily log
- 156: all

Batteries
- Continue charging, testing, inventory
6/7/2014 158 - H4 BDA 20s spikes
- H4 hf ramp burst: 12:05, 15:20, 20;40
2.6394 - elec noise: H4
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.709 99.409 99.781 99.152  
6/8/2014 159 - H4 BDA 20s spikes
- H4 ramp burst: 06:40
- L2 hf ramp burst: 07:30
- H4 hf ramp burst: 11:00, 15:20, 19:50
2.1431 - elec noise: H4
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all except H2 clk diff too large
- 2 frames from I57 to IDC delayed
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.710 99.413 99.782 99.157  
6/9/2014 160 - H4 BDA 20s spikes
- H4 ramp burst: 07:45
- H4 hf ramp burst: 03:20, 13:00, 14:25, 19:15
- L2 hf ramp burst: 03:50, 09:55
2.1899 - elec noise: H4, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.712 99.416 99.784 99.163   Daily log
- Reprocess: 158-159
- All: 157-159

Batteries
- Continue testing/charging/inventory

Michael status call
- Discuss otay/pinon trip
6/10/2014 161 - H4 BDA 20s spikes
- L1 single spike: 07:35
- H4 ramp burst: 08:20
- L2 ramp burst: 10:30
- H4 hf ramp burst: 03:40, 19:50
- L2 hf ramp burst: 04:00, 15:05
1.7597 - 16:33 (0.1pa 115) - elec noise: H4, L2
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 3 frames from profile2 to UCSD delayed
- 2 frames from profile2 to UCSD duplicated
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.714 99.420 99.785 99.168  
6/11/2014 162 - H4 BDA 20s spikes 2.6646 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.716 99.424 99.786 99.173   Daily log
- Reprocess: 161
- All: 160-161

Batteries
- Continue charging/labeling, etc…

Randy budget question
- Discuss with michael
- Call randy
- Ask lenore for #’s

SSI
- Download and file 32bit and 64bit versions from sandia
6/12/2014 163 - H4 BDA 20s spikes
- H4 hf ramp burst: 05:30
2.4164 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.718 99.427 99.788 99.178  
6/13/2014 164 - AS108 revalidation tests - H4 BDA 20s spikes
- H4 hf ramp burst: 05:45
2.29 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 6 frames from I57 to IDC delayed
- 4 frames from I57 to IDC duplicated
- 2 frames from I57 to NDC duplicated
- 2 frames from I57 to UCSD duplicated
- site visit related flags
  100.000 100.000 99.977 100.000 x 100.000 100.000 99.998 100.000 x 99.719 99.431 99.789 99.183  
6/14/2014 165 - H4 BDA 20s spikes
- H4 hf ramp burst: 22:10
2.4102 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- I57H2 BDF flag
  100.000 100.000 100.000 100.000 x 100.000 100.000 99.998 100.000 x 99.721 99.434 99.790 99.188  
6/15/2014 166 - H4 BDA 20s spikes
- H4 hf ramp burst: 05:00
2.5742 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all except H3 clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 99.998 100.000 x 99.723 99.437 99.791 99.193  
6/16/2014 167 - NTR 2.4508 - 04:40 (0.2pa 002) - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 99.999 100.000 x 99.724 99.441 99.793 99.198   Daily log
- Reprocess: 164 - 166
6/17/2014 168 - install temp as108 display computer, fix links and start large backfill - NTR 2.6848 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all except H3 clk diff too large
- 449 frames from I57 to IDC delayed
- 15 frames from I57 to IDC duplicated
- 837 frames from I57 to NDC delayed
- 24 frames from I57 to NDC duplicated
- 4 frames from I57 to UCSD duplicated
- site visit relayed flags
  99.931 100.000 99.491 98.750 x 99.996 100.000 99.969 99.926 x 99.726 99.444 99.791 99.195  
6/18/2014 169 - H4 ramp burst: 12:30
- H4 hf ramp burst: 15:00
2.1492 - 20:56 (0.1pa 292) - elec noise: H4
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.996 100.000 99.970 99.931 x 99.727 99.447 99.792 99.200   Daily log
- Reprocess: 167-168

Batteries
- Take home to continue testing
- Submit or
6/19/2014 170 - install chavez as108 calibration computer - H4 hf ramp burst: 10:05, 15:45 1.7271 - 17:29 (0.1pa 111) - elec noise: H4
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- site visit related flags
  100.000 100.000 100.000 100.000 x 99.996 100.000 99.972 99.934 x 99.729 99.451 99.793 99.205  
6/20/2014 171 - H4 ramp burst: 04:40
- H4 hf ramp burst: 00:05, 13:00, 14:30, 21:50
- L2 hf ramp burst: 05:00
1.7651 - 21:13 (0.1pa 046)
- 21:50
(0.1pa 094)
- elec noise: H4
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 4 frames from profile2 to UCSD duplicated
  100.000 100.000 100.000 100.000 x 99.997 100.000 99.973 99.938 x 99.730 99.454 99.795 99.209   OR - 103551 Daily log
- Reprocess: 170

Batteries
- Continue testing/charging from home
- Update OR with requested details

Trip return
- Write reports

Work on timesheet with michael
L4
- Inspect battery enclosure in preparation for upgrade
6/21/2014 172 - H4 hf ramp burst: 06:10, 19:30
- L2 hf ramp burst: 16:10
1.9469 - elec noise: L3
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.997 100.000 99.975 99.940 x 99.732 99.457 99.796 99.214  
6/22/2014 173 - NTR 2.5408 - elec noise:
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.997 100.000 99.976 99.943 x 99.733 99.460 99.797 99.218  
6/23/2014 174 - H4 hf ramp burst: 20:00 2.1352 - elec noise:
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all except L3 clk diff too large   100.000 100.000 100.000 100.000 x 99.997 100.000 99.977 99.946 x 99.735 99.463 99.798 99.223   Daily log
- Reprocess: 171-173

Randy – wnrs
- Discuss on phone, give another drawing

Michael
– status call

Timesheet
- complete, copy, submit

Trip preparation
- Gather equipment for tomorrow
6/24/2014 175 - install new batteries at L3 - H4 hf ramp burst: 07:05, 20:05 - N/A - all except H3 clk diff too large
- site visit related flags
  99.479 99.690 100.000 99.479 x 99.975 99.987 99.978 99.926 x 99.734 99.465 99.799 99.224  
6/25/2014 176 - H4 hf ramp burst: 07:10, 19:10 2.018 - elec noise:
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 1120 frames from I57 to NDC missing
  100.000 100.000 100.000 74.236 x 99.976 99.988 99.979 98.899 x 99.735 99.468 99.801 99.082   Batteries
- update inventory
- continue charging/testing from home
6/26/2014 177 - H4 hf ramp burst: 08:10,
- L2 hf ramp burst: 07:30
2.1264 - elec noise:
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 4320 frames from I57 to NDC mission
  100.000 100.000 100.000 0.000 x 99.977 99.988 99.980 95.095 x 99.737 99.471 99.802 98.523  
6/27/2014 178 - H4 hf ramp burst: 20:05 2.4339 - elec noise: L3
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- I57L2 BDF flag
- I57H1 BDF flag
- I57H2 BDF flag
- 8458 frames from I57 to NDC delayed
- 5430 frames from I57 to NDC backfilled
  100.000 100.000 100.000 100.000 x 99.978 99.989 99.980 95.276 x 99.738 99.474 99.803 98.531  
6/28/2014 179 - H4 hf ramp burst: 06:50, 19;40 2.1613 - elec noise:
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all except H2 clk diff too large   100.000 100.000 100.000 100.000 x 99.979 99.989 99.981 95.445 x 99.740 99.477 99.804 98.539  
6/29/2014 180 - H4 ramp burst: 08:20
- L2 hf ramp burst: 08:25
- H4 hf ramp burst: 18:50
1.8337 - elec noise: L2
- wnrs noise: H1, L1-L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.980 99.989 99.982 95.602 x 99.741 99.479 99.805 98.547  
6/30/2014 181 - H4 hf ramp burst: 09:20
- L2 hf ramp burst: 09:40, 15:25
2.054 - elec noise: L2
- wnrs noise: L1-L2
- 8hz bump: L1-L4
- all clk diff too large
- 1 frame from I57 to IDC delayed
  100.000 100.000 100.000 100.000 x 99.980 99.990 99.982 95.749 x 99.742 99.482 99.806 98.555   Daily log
- Uptime – catchup

Batteries
- Take next set home and start charging/testing

Inventory
- Ask delia for more information
                    x x x x x x x x x x x x x x          
7/1/2014 182 - all except H2 clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.744 99.485 99.807 98.563   Daily log
- Uptime

Batteries
- Continue charging/testing
7/2/2014 183 - Vandenberg Delta II launch at 09:56 - all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.745 99.488 99.808 98.571   MR - 103785 Daily log
- Uptimes

Batteries
- Continue charging/testing

MR
- File sent emails
- Submit report

Inventory
- Search for delia digitziers but could not find
- Respond saying I can’t figure out what they are referring to.
7/3/2014 184 - all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.747 99.491 99.809 98.579  
7/4/2014 185 - all clk diff too large
- 1 frame from profile2 to UCSD delayed
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.748 99.494 99.810 98.586  
7/5/2014 186 - all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.749 99.496 99.811 98.594  
7/6/2014 187 - all except H2 clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.751 99.499 99.812 98.602  
7/7/2014 188 - all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.752 99.502 99.813 98.609   Daily log
- Uptimes

Batteries
- Continue testing/charging
- Perform analysis on L3 swap, found that new batteries drop 3% less each night than old

Michael status
- Have meeting – discuss all

Randy grounding
- Look through past notes, give values and procedures
7/8/2014 189 - all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.753 99.504 99.814 98.616   Daily log
- Uptimes

Batteries
- Continue testing/charging
- Ask michael if I can purchase 2 more chargers
- Record results in inventory
- Make labels
- Purchase 2 additional battery chargers

Tantor
- Update scripts to include new, daily battery voltage drop %’s in the website year trends

Prep for tomorrow’s trip
- Change laptop settings
- Pack equipment
- Fuel truck

Signal analysis
- Ask michael if we can give Stephen a login so he can get data
- Ask netops for Stephen login
7/9/2014 190 - install new batteries at L4
- AS108 revalidation tests
- all except H2, H3 clk diff too large
- site visit related flags
  99.579 99.750 100.000 99.579 x 99.953 99.972 100.000 99.953 x 99.752 99.506 99.815 98.621   L4
- Replace Batteries
7/10/2014 191 - all clk diff too large   100.000 100.000 100.000 100.000 x 99.958 99.975 100.000 99.958 x 99.754 99.508 99.816 98.629   Daily log
- Uptimes

Trip return
- Unload

Battery swap
- Start charging old L3/L4 batteries
- Send PR for yesterday’s work

Arrowsmith
- Discuss giving him scp access to archive with netops
7/11/2014 192 - all clk diff too large   100.000 100.000 100.000 100.000 x 99.962 99.977 100.000 99.962 x 99.755 99.511 99.817 98.636   PR - 104166 Daily log
- Uptimes

Arrowsmith
- Test netops login for Stephen

Batteries
- Receive 2 new chargers (delivered to wrong location)
- Submit express card paperwork
- Start charging all 4 batteries simultaneously
7/12/2014 193 - all clk diff too large   100.000 100.000 100.000 100.000 x 99.965 99.979 100.000 99.965 x 99.756 99.513 99.818 98.643  
7/13/2014 194 - all clk diff too large   100.000 100.000 100.000 100.000 x 99.968 99.981 100.000 99.968 x 99.757 99.516 99.819 98.650  
7/14/2014 195 - all except L3 clk diff too large   100.000 100.000 100.000 100.000 x 99.970 99.982 100.000 99.970 x 99.759 99.518 99.820 98.657   CCN - 104235 Daily log
- Uptimes
- Reprocess: 192-194

Michael status call:
- discuss all

Submit ccn
7/15/2014 196 - all clk diff too large
- 193 frames from I57 to NDC delayed
  100.000 100.000 100.000 100.000 x 99.972 99.983 100.000 99.972 x 99.760 99.521 99.821 98.664   Daily log
- Uptimes
7/16/2014 197 - all clk diff too large
- 1 frame from I57 to IDC delayed
  100.000 100.000 100.000 100.000 x 99.974 99.984 100.000 99.974 x 99.761 99.523 99.822 98.670   Daily log
- Uptimes

Ndc outage
- Informed michael that the IDC alert occurred while backfilling ndc
7/17/2014 198 - all clk diff too large
- 1 frame from profile2 to UCSD duplicated
  100.000 100.000 100.000 100.000 x 99.975 99.985 100.000 99.975 x 99.762 99.526 99.823 98.677   Daily log
- Uptimes
- All: 161

Batteries
- Test 4 of the batteries returned from L3/L4
• All are bad – about 13ah
7/18/2014 199 - all clk diff too large
- 1 frame from profile2 to UCSD delayed
  100.000 100.000 100.000 100.000 x 99.977 99.986 100.000 99.977 x 99.764 99.528 99.824 98.684  
7/19/2014 200 - all clk diff too large   100.000 100.000 100.000 100.000 x 99.978 99.987 100.000 99.978 x 99.765 99.530 99.825 98.690  
7/20/2014 201 - all clk diff too large   100.000 100.000 100.000 100.000 x 99.979 99.987 100.000 99.979 x 99.766 99.533 99.826 98.697  
7/21/2014 202 - all except H2 clk diff too large
- 27 frames from I57 to NDC delayed
  100.000 100.000 100.000 100.000 x 99.980 99.988 100.000 99.980 x 99.767 99.535 99.826 98.703  
7/22/2014 203 - all except H3 clk diff too large   100.000 100.000 100.000 100.000 x 99.981 99.989 100.000 99.981 x 99.768 99.537 99.827 98.710   Daily log
- Uptimes

Randy
- Return call, wants me to overnight keller
- Package and ship keller
7/23/2014 204 - all clk diff too large
- 6 frames from I57 to NDC delayed
  100.000 100.000 100.000 100.000 x 99.982 99.989 100.000 99.982 x 99.769 99.539 99.828 98.716   Daily log
- Uptimes
- Reprocess 203

Michael status
– quick email with update

Batteries
- Continue testing old L3/L4 batteries
- Start writing script to analyze L3 before and after exchange
- Long email to michael/randy with all stats regarding exchange

Clean lab

Trip prep
- Fuel truck, load equip
7/24/2014 205 - site inspection - all clk diff too large
- BWS flag
- site visit related flags
  100.000 100.000 100.000 100.000 x 99.982 99.990 100.000 99.982 x 99.770 99.542 99.829 98.722   All - inspection
- Drove by all sites, performed exterior inspection
- L1 – Many inlets buried
- H1 – no inlets buried
- Many fallen trees throughout compound, no fence or filter damage
- Interior vault inspection
- L1, L2, H4

CRF
- Power cycle UPS USB
- Still complaining about battery replacement
- Status shows current battery installed on 11.19.08 and has 31 min runtime
7/25/2014 206 - all clk diff too large   100.000 100.000 100.000 100.000 x 99.983 99.990 100.000 99.983 x 99.772 99.544 99.830 98.728  
7/26/2014 207 - all clk diff too large   100.000 100.000 100.000 100.000 x 99.984 99.990 100.000 99.984 x 99.773 99.546 99.831 98.735  
7/27/2014 208 - all clk diff too large   100.000 100.000 100.000 100.000 x 99.984 99.991 100.000 99.984 x 99.774 99.548 99.832 98.741  
7/28/2014 209 - all clk diff too large
- 22 frames from I57 to NDC delayed
  100.000 100.000 100.000 100.000 x 99.985 99.991 100.000 99.985 x 99.775 99.551 99.832 98.747   Batteries
- Research new price quote
- Send quote to michael/curt
- Continue testing old L3/L4 batteries

Randy call
- Wind filter question
- Discuss budget stuff –
• He will ask curt about ssi costs
• They will include costs for pipes and 9th element and element move
• We aren’t going to worry about vault upgrade now
• They have my battery costs

Write 7/24 trip reports

Funding
- Spend several hours researching costs for many lab items and provide detailed summary to michael
- Additional research on apple computers for customization on Michael’s system
7/29/2014 210 - all clk diff too large
- 838 frames from I57 to NDC missing
  100.000 100.000 100.000 80.486 x 99.985 99.991 100.000 99.313 x 99.776 99.553 99.833 98.660   Daily log
- Uptimes
7/30/2014 211 - all except H2 clk diff too large
- 3100 frames from I57 to NDC delayed
- 838 frames from I57 to NDC backfilled
  100.000 100.000 100.000 100.000 x 99.986 99.992 100.000 99.335 x 99.777 99.555 99.834 98.666   Daily log
- Uptimes

Randy
- Answer question about key_id
7/31/2014 212 - all clk diff too large   100.000 100.000 100.000 100.000 x 99.986 99.992 100.000 99.357 x 99.778 99.557 99.835 98.672   Daily log
- Uptimes

Trip prep
- Fuel truck
- Load supplies
                    x x x x x x x x x x x x x x          
8/1/2014 213 - convert squ 64/70 to 64/35 - all clk diff too large
- site visit related flags
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.779 99.559 99.836 98.679   H2
- The temperature label on the digitizer claims that 100 was reached but not 125
8/2/2014 214 - power outage caused system to go offline - all except L3 clk diff too large
- 1 frame from I57 to IDC delayed
- 1131 frames from I57 to IDC missing
- 1132 frames from I57 to NDC missing
- 1132 frames from I57 to UCSD missing
- 1126 frames from profile2 to UCSD missing
  73.750 73.889 73.727 73.750 x 86.875 86.944 86.864 86.875 x 99.657 99.439 99.712 98.562  
8/3/2014 215 - manually restart software processes N/A   40.694 0.000 40.556 40.694 x 71.481 57.963 71.428 71.481 x 99.383 98.976 99.435 98.293  
8/4/2014 216 - 2720 frames from profile2 to UCSD missing
- outage related flags
  100.000 37.083 100.000 100.000 x 78.611 52.743 78.571 78.611 x 99.386 98.690 99.437 98.301   PR - 104629 Daily log
- Reprocess: 214, 215
- Uptimes: 204

Weekend outage
- Give michael brief update on suspected cause
- 15:05 – restart profile2 software processes’
- Outage details:
• 17:42 – current levels spiked and then went offline
• 18:02 – log shows came back online but could not connect
• UPS reports 3 power outages:
• 8/2 – 17:42 (24 seconds)
• 8/2 – 19:10 (3 seconds)
• 8/3 – 14:20 (18 seconds)
- According to UPS power failed at 17:41:42 and the UPS shutdown at 17:42:06
- Power came back online it seems and the UPS was suffieciently charged to turn on at 18:02
- Forwarder logs stop at 18:04 for unknown reason, server continues to attempt connections after that until I restart software next day
- Respond to PR
- Forward PR to U.S. with additional explanation

Michael
- Email regarding message system, this week’s schedule and next week’s vacation
- Email regarding temperature strips
8/5/2014 217 - enable frame verification on umoya ndc box and restarts at 17:38, 17:53 - receiver restart related flags   100.000 100.000 100.000 100.000 x 82.889 62.194 82.857 82.889 x 99.389 98.696 99.440 98.309   Authentication
- Discuss Hawaii problems with randy on phone
- Download certificates from web portal
- Upload certificates to umoya
- Modify cdtools.par and restart receiver twice
- Analyze logs, unable to determine status of umoya authentication checks
- Inform randy won’t have more time until Thursday to figure it out
8/6/2014 218 - all except H2 clk diff too large
- 9 frames from profile2 to UCSD delayed
- 4 frames from profile2 to UCSD duplicated
  100.000 100.000 100.000 100.000 x 85.741 68.495 85.714 85.741 x 99.392 98.702 99.442 98.316  
8/7/2014 219 - all clk diff too large
- 4 frames from profile2 to UCSD delayed
- 4 frames from profile2 to UCSD duplicated
  100.000 100.000 100.000 100.000 x 87.778 72.996 87.755 87.778 x 99.394 98.708 99.445 98.324   Daily log
- Uptimes: 205 – 218

Write all trip reports

Timesheet
– complete, copy, submit

Authentication
- Discuss met key id with randy
- Send him excerpt from log

Express card
- Submit home depot receipts

Website
- Remove login requirement
- Create new script that fills in bad days like day 215

Soh script
- Modify to send daily summary text
8/8/2014 220 - all clk diff too large
- 4 frames from profile2 to UCSD duplicated
  100.000 100.000 100.000 100.000 x 89.306 76.372 89.285 89.306 x 99.397 98.714 99.448 98.332   MR - 104831 Daily log
- Uptimes

Submit MR

Batteries
- Continue testing l3/l4 batteries
8/9/2014 221 - all except H2, H3 clk diff too large
- 4 frames from profile2 to UCSD delayed
- 2 frames from profile2 to UCSD duplicated
  100.000 100.000 100.000 100.000 x 90.494 78.997 90.476 90.494 x 99.400 98.720 99.450 98.339  
8/10/2014 222 - all clk diff too large
- 4 frames from I57 to IDC delayed
- 1 frame from I57 to IDC duplicated
- 1 frame from I57 to NDC duplicated
- 1 frame from I57 to UCSD duplicated
- 3 frames from profile2 to UCSD delayed
- 1 frame from profile2 to UCSD duplicated
  100.000 100.000 100.000 100.000 x 91.444 81.097 91.428 91.444 x 99.403 98.725 99.453 98.347  
8/11/2014 223 - all clk diff too large
- 2 frames from I57 to IDC delayed
- 1 frame from profile2 to UCSD delayed
- 2 frames from profile2 to UCSD duplicated
  100.000 100.000 100.000 100.000 x 92.222 82.816 92.208 92.222 x 99.405 98.731 99.455 98.354   Daily log
- Uptimes: 219-222
- Reprocess: 221-222
8/12/2014 224 - all except L3 clk diff too large
- BWS error flag
  100.000 100.000 100.000 100.000 x 92.870 84.248 92.857 92.870 x 99.408 98.737 99.458 98.362  
8/13/2014 225 - all except H1, H3 clk diff too large   100.000 100.000 100.000 100.000 x 93.419 85.459 93.406 93.419 x 99.411 98.742 99.460 98.369   Daily log
- Uptimes: 223-224

Express card
- Pickup new card from lenore

Batteries
- Give jay info on battery purchase and explain we don’t need spares

Wgb annual summary
- Provide summary to michael with events at I57 for last year
- Work on phone with michael to write curt response

Jay – pki
- Inform that I can upload keys on Monday after testing on web portal
8/14/2014 226 - all clk diff too large
- 34 frames from I57 to IDC delayed
- 1 frame from I57 to IDC duplicated
- 207 frames from I57 to NDC delayed
- 1 frame from I57 to NDC duplicated
- 66 frames from I57 to NDC missing
- 213 frames from I57 to UCSD delayed
- 1 frame from I57 to UCSD duplicated
- 106 frames from profile2 to UCSD delayed
- 3 frames from profile2 to UCSD duplicated
  100.000 100.000 100.000 98.500 x 93.889 86.498 93.877 93.782 x 99.413 98.748 99.462 98.369  
8/15/2014 227 - all clk diff too large
- 1413 frames from I57 to NDC delayed
- 66 frames from I57 to NDC backfilled
- 1263 frames from I57 to UCSD delayed
- 1296 frames from profile2 to UCSD delayed
- 7 frames from profile2 to UCSD duplicated
  100.000 100.000 100.000 100.000 x 94.296 87.398 94.286 94.196 x 99.416 98.753 99.465 98.377  
8/16/2014 228 - somebody may have tried to open door but hasp stopped them? - all clk diff too large
- 1 frame from profile2 to UCSD delayed
- vault door open flag
  100.000 100.000 100.000 100.000 x 94.653 88.186 94.643 94.559 x 99.418 98.759 99.467 98.384  
8/17/2014 229 - all clk diff too large   100.000 100.000 100.000 100.000 x 94.967 88.881 94.958 94.879 x 99.421 98.764 99.470 98.391  
8/18/2014 230 - all clk diff too large   100.000 100.000 100.000 100.000 x 95.247 89.498 95.238 95.164 x 99.423 98.770 99.472 98.398   Daily log
- Uptimes: 225 -229

Crf door
- It looks like somebody tried to open the crf door while it was still locked at 5pm on Saturday, both attempts were less than a second
- Data confirms that power or temperature did not change
- Email michael with details

Michael status call
- Write email to ask curt about UPS
- Discuss crf door
- Discuss funding
- He would like additional information in the alert text (how many frames are missing)

UPS
- Provide UAF with quote on replacement battery and network card

Wnrs
- Analyze data before and after the latest squ upgrade:
• The new smaller size eliminates the resonance without the use of capillaries
• It does not provide as good of wind filtering though
- Need to test 4 pods with 4 sensors and post collection digital averaging
- Put together cad drawings to estimate size and phase response, confirm that straight averaging with no phase correction should work
- Ask michael if I can proceed
8/19/2014 231 - enter CRF to retrieve some equipment - all except H2 clk diff too large
- 5 frames from profile2 to UCSD delayed
- 2 frames from p rofile2 to UCSD duplicated
- site visit related flags
  100.000 100.000 100.000 100.000 x 95.497 90.051 95.489 95.418 x 99.426 98.775 99.474 98.405   Daily log
- Uptimes
8/20/2014 232 - move all equipment from UPS to wall power, move PC's to as108 ups - storms caused power outage, then UPS failure, then baytech failure - N/A   76.218 98.889 76.821 76.218 x 94.533 90.493 94.555 94.458 x 99.326 98.776 99.376 98.309   PR - 105089 57 power outage
- Went offline at about 5:20
- Profile2 came back online
- Cannot ping I57
- Cannot communicate with baytech

Submit PR
CRF
- Upon arrival found the UPS on but complaining about battery replacement
- Also found that both the primary and backup computers were powered up
- This was causing a network conflict
- The baytech did not show any soh lights or anything on the lcd. It was in failure mode which provides power to all outlets by default
- Attempted to “reset” UPS outlet by power cycling UPS
- The UPS complained loudly and switched harshly
- Decided the UPS was no longer functioning correctly and took offline
- The UPS cycle did not help the baytech but the UPS outlet was an unknown and we do not have any other 20a outlets at pinon to test the baytech with
- Remove the baytech from the system and bring back to lab for testing (later confirmed it was damaged beyond repair)
- Used one power strip plugged directly into wall to power everything that was previously on ups
- Use second power strip plugged into the as108 ups to power the computers
- Powered down backup computer, brought primary online
8/21/2014 233 - another power outage took system offline until software could be fixed and restarted
- when brought back online BWD was out of range after that?
- all clk diff too large
- all sites gps rcv off?
- 52 frames from I57 to IDC missing
- 52 frames from I57 to NDC missing
- 52 frames from I57 to UCSD missing
- 140 frames from profile2 to UCSD missing
- BWD out of range
  98.796 96.759 98.796 98.796 x 94.736 90.791 94.757 94.665 x 99.324 98.767 99.373 98.311   UPS
- Give michael scenarios on waiting a few months to replace equipment

Power controller
- Research – find unit and pricing – provide to michael

Write trip report

Daily log
- Uptimes
- Reprocess: 220

Antelope
- Research license – unclear if need to be renewed or not
- Attempt command from hal – unsuccessfull
- Email michael asking if he renews his each year

Michael – status call
- Purchase power strip and ups accessories now
- Inform UAF that we purchased

Order power strip and ups supplies

Mac
- Research osx updates – request assistance from help desk
- Update terminal settings
- Update bashrc settings
- Update nfs automount settings

57 outage
- 20:40 - server froze while performing rollover for some reason?
- 20:52 – restart software
- 20:57 – manually start forwarder after restarts lost sync with server
- 21:20 – restart profile2 software
• Complaining about missing file error
- 21:25 – cleanup tables, clear queues, restart again
- Email michael
8/22/2014 234 - BWD malfunction - all clk diff too large
- BWD out of range
  100.000 100.000 100.000 100.000 x 94.975 91.210 94.995 94.907 x 99.327 98.772 99.376 98.318  
8/23/2014 235 - BWD malfunction - all clk diff too large
- BWD out of range
  100.000 100.000 100.000 100.000 x 95.194 91.592 95.213 95.129 x 99.329 98.777 99.379 98.326  
8/24/2014 236 - BWD malfunction - all clk diff too large
- BWD out of range
  100.000 100.000 100.000 100.000 x 95.394 91.943 95.413 95.332 x 99.332 98.783 99.381 98.333  
8/25/2014 237 - BWD malfunction - all except H2 clk diff too large
- 1 frame from profile2 to UCSD delayed
- BWD out of range
  100.000 100.000 100.000 100.000 x 95.578 92.265 95.596 95.518 x 99.335 98.788 99.384 98.340   PR - 105177 Daily log
- Uptimes: 233:236

Ssi computer
- Put together new quote for fanless mini pc and amazon accessories and send to UAF

Michael status call
- Find out when UPS equipment will arrive
- Delay flim flom until after UPS install
- Get quote for truck roof
- Training for michael to run 57 in my absence

Website
- Update publications page with michael & catherines requests
- Ask them to confirm everything looks correct

BWD problem
- Started during last power outage
- Goes negative
- Submit PR
8/26/2014 238 - BWD malfunction - all clk diff too large
- BWD out of range
  100.000 100.000 100.000 100.000 x 95.748 92.562 95.765 95.691 x 99.338 98.793 99.386 98.347   Daily log
- Uptimes: 237

BWD malfunction
- Print out circuit diagram with troubleshooting points to check
- Print out past PR with same problem to follow

Power strip
- Unpack – submit paperwork
- Establish connection, change IP, SMTP, login settings for both units
- Update network and power diagrams

UPS
- Receive – unpack
8/27/2014 239 - BWD malfunction - all clk diff too large
- BWD out of range
  100.000 100.000 100.000 100.000 x 95.906 92.838 95.922 95.850 x 99.341 98.798 99.389 98.354   Daily log
- Uptimes: 238
- All: xxx - 238

Mac
- Ask paul if he can perform OS update tomorrow

Make final arrangements
- Load supplies, fuel truck
8/28/2014 240 - UPS battery replacement
- install remote managed power strips
- conduct power outage test
- BWD malfunction - all clk diff too large
- 182 frames from I57 to IDC missing
- 182 frames from I57 to NDC missing
- 182 frames from I57 to UCSD missing
- 76 frames from profile2 to UCSD missing
- site visit related flags
- BWD out of range
  95.743 98.241 95.625 95.743 x 95.900 93.031 95.912 95.846 x 99.326 98.795 99.373 98.343   Mac OS updated by netops CRF
• Power down all
• Install new power controller strips on back of rack
• Install new UPS battery
• Power all up and confirm working
• Configure power strips with items plugged in
• Label power cords
• Attempt to configure UPS network card, unable to get it to communicate with UPS
• Perform power outage test, UPS handled load fine until generator supplied power
8/29/2014 241 - BWD malfunction - all except H1 clk diff too large
- BWS high winds
- BWD out of range
  100.000 100.000 100.000 100.000 x 96.041 93.271 96.053 95.990 x 99.328 98.800 99.376 98.350   Daily log
- Uptimes

Tantor
- Attempt to fix after OS update – unsuccessfull

Provide fee with shipping address for ssi pc
8/30/2014 242 - BWD malfunction - all clk diff too large
- BWD out of range
  100.000 100.000 100.000 100.000 x 96.173 93.495 96.184 96.123 x 99.331 98.805 99.378 98.356   PR - 105089 Submit PR
8/31/2014 243 - BWD malfunction - all except L3 clk diff too large
- L1, L2, L3, H3, H4 BDF out of range
- L1 BWD out of range
  100.000 100.000 100.000 100.000 x 96.297 93.705 96.307 96.248 x 99.334 98.810 99.381 98.363  
                    x x x x x x x x x x x x x x x        
9/1/2014 244 - BWD malfunction - H4 BDA 20s spikes
-
H4 hf ramp burst: 06;40, 20:20
- L2 hf ramp burst: 10:30
2.5907 - elec noise:
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- L1 BWD out of range
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.337 98.815 99.384 98.370  
9/2/2014 245 - BWD malfunction - H4 BDA 20s spikes
- H4 ramp burst: 06:30
- L2 hf ramp burst: 08:15
- H2 hf ramp burst: 14;20
- L1 spiky noise all day
2.1017 - elec noise: L1, H2
- wnrs noise: L1-L2
- 8hz bump: L1-L4
- all clk diff too large
- L1 BWD out of range
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.339 98.820 99.386 98.377   Daily log
- Uptimes: 241-244
- Reprocess:
- Update with functions performed last week

Tantor
- Continue post OS update repairs

Timesheet
- Give michael numbers in email for verification
- Fill out and submit

Write trip report for UPS install

Michael status call
- Discuss pinon, timesheet, Canada trip, ssi, etc…
9/3/2014 246 - BWD malfunction - H4 BDA 20s spikes
- H4 hf ramp burst: 06:25, 20:35
- L1 spiky noise all day
2.3497 - elec noise: L3
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all except L3 clk diff too large
- L1 BWD out of range
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.342 98.825 99.389 98.383   Daily log
- Uptimes: 245
- Reprocess:239:244

Tantor
- Continue post OS update repairs
- Determine it’s a java problem that shutsdown matlab when a 2nd window is opened
- Install java
• Still not working
- Update matlab
• Script is working now!

Antelope
- Attempt to register again, perl error
- Ask Geoff davis for assistance – he showed me how to source the environment
- I was then able to run the registration program which confirmed I sent one off to brtt

Reports
- File emails for MR
9/4/2014 247 - BWD malfunction - H4 BDA 20s spikes
- H4 ramp burst: 06:10
- L2 hf ramp burst: 06:55, 16:35
- H4 hf ramp burst: 21:05
- L1 spiky noise all day
1.8978 - elec noise: L1, H4
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- L1 BWD out of range
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.345 98.830 99.391 98.390   PR - 105177
MR - 105442
Daily log
- Uptimes: 246
- Reprocess: 244-246

Randy call
- Device to power cycle the bwd circuit
- Using stealth computers throughout the network
- Reporting
- New guys come out to pinon

Clean
- File paperwork
- Unload truck
- Organize tools

Pinon trip prep
- Make notes
- Gather materials
- Respond to reports

Batteries
- Test last of l3/l4 batteries
- Update inventory with all of the L3/L4 battery information

• Reports
- Submit August MR
- Submit UPS battery ccn
- Submit BWD PR update
9/5/2014 248 - BWD malfunction - H4 BDA 20s spikes
- H4 ramp burst: 05:20
- L2 hf ramp burst: 06:05
- H1 hf ramp burst: 12:50, 16:20
- H2 hf ramp burst: 12:15, 16:15
- L1 spiky noise all day
1.7184 - 16:50 (0.1pa 105) - elec noise: H1, H2, H4, L1
- wnrs noise: L1-L3
- 8hz bump: L1-l4
- all clk diff too large
- L1 BWD out of range
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.347 98.834 99.394 98.396   Daily log
- All: 247

Inventory
- Respond to megan about sun work station
- Pull out parche and thorpe and update inventory

Continue cleaning lab
- Work with megan to get surplus pickup scheduled
• take pics, give details
- work with megan again to schedule hazmat pickup
• sort batteries and put on pallet
• take pics, give details

express card
- give lenore invoice on ups order
- ask lenore how to get mcmaster invoice she wants
9/6/2014 249 - 22:10 - CRF AC offline
- BWD malfunction
- H4 BDA 20s spikes
- H4 ramp burst: 04:15
- L2 hf ramp burst: 06:00
- H2 hf ramp burst: 10:30, 16:00
- H1 hf ramp burst: 12:10, 15:25
- H4 hf ramp burst: 22:10
- L1 spiky noise all day
1.3213 - 16:42 (0.6pa 105) - elec noise: H1, H2, H4, L1
- wnrs noise: L1-L3
- 8hz bump: L1-l4
- all clk diff too large
- 191 frames from I57 to NDC missing
- L1 BWD out of range
  100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.350 98.839 99.396 98.396  
9/7/2014 250 - CRF AC offline
- BWD malfunction
- H4 BDA 20s spikes
- H4 ramp burst: 03:15
- L2 hf ramp burst: 06:45
- H1 hf ramp burst: 10:40, 15;40
- H2 hf ramp burst: 09:30, 15;30
- L1 spiky noise all day
1.1643 - elec noise: H4, L1
- wnrs noise: L1-L2
- 8hz bump: L1-L4
- all clk diff too large
- H1 GPS unlocked
- 191 frames from I57 to NDC backfilled
- CF VTP out of range
- L1 BWD out of range
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.353 98.844 99.398 98.403  
9/8/2014 251 - CRF AC offline
- BWD malfunction
- H4 BDA 20s spikes
- H1 ramp burst: 09:00
- H2 hf ramp burst: 04:50, 20:35
- L2 hf ramp burst: 06:20
- H1 hf ramp burst: 17:55
- L1 spiky noise all day
1.0263 - elec noise: L1, L2, H1, H2
- wnrs noise: L2
- 8hz bump: L1-L4
- all except H1 clk diff too large
- CF VTP out of range
- L1 BWD out of range
  100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.355 98.848 99.401 98.403   Daily log
- Uptimes: 248-250
- Reprocess: 248-250

Antelope
- License expired which is why it didn’t work over weekend
- Re-install per meyer’s direction

Michael status
- Discuss timesheet
- Discuss truck and whether to use fleet services or do on our own

CRF AC
- Went offline on 248-22:10 at the same time the CRF experienced a power surge.
- Assume the breaker is tripped
- The CRF is staying at about 39c but is peaking as high as 41
- Ask but nobody is heading to pfo tomorrow
- Inform michael
9/9/2014 252 - CRF AC offline
- BWD malfunction
- H4 BDA 20s spikes
- H1 hf ramp burst: 07:00, 16:45
- H2 hf ramp burst: 05:00, 17:40
- L2 hf ramp burst: 05:40
- L1 spiky noise all day
1.3786 - elec noise: H1, L1
- wnrs noise: L2, L3
- 8hz bump: L1-L4
- all except H2 clk diff too large
- 93 frames from I57 to IDC delayed
- CF VTP out of range
- L1 BWD out of range
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.358 98.853 99.403 98.409   Daily log
- All: 251

Trip preparation
- Load/fuel truck
- Clean, fuel, test chainsaw
- Create work plan for tomorrow

Crf temp
- Monitor throughout day, did not go over 41
9/10/2014 253 - reset CRF AC breaker
- reset BWS, BWD channels
- CRF AC offline
- BWD malfunction
- H4 BDA 20s spikes
- H4 ramp burst: 03:10
- H1 ramp burst: 07:35
- H4 hf ramp burst: 00:10
- H2 hf ramp burst: 06:05
- L2 hf ramp burst: 05:30
- H1 hf ramp burst: 17:20
- L1 spiky noise until BWD repair
1.0049 - 18:33 (0.2pa 338) - elec noise: H1, H2, L1
- wnrs noise: L2
- 8hz bump: L1-L4
- all clk diff too large
- CF VTP out of range
- L1 BWD out of range
- Site visit related flags
  100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.360 98.857 99.406 98.415   CRF
• Reset Air Con Breaker

L1
• Power cycle wind sensor did not fix BWD offset problem
• Physically disconnect BWD output from smart24
-This fixed problem but I later found out it transferred the issue to the BWS channel
• Physically disconnect BWS output from smart24
-This fixed problem again
• There must be some sort of buildup between the backend of the opamp and the digitizer that can only be released with a physically disconnect? One would think that power cycling the op-amp would relieve this but it does not.

H1
• Use chainsaw to cut up and remove tree blocking access to the H1 entrance
9/11/2014 254 - H4 BDA 20s spikes
- H4 ramp burst: 02:50
- H1 ramp burst: 06:40
- H4 hf ramp burst: 00:45, 23:40
- H2 hf ramp burst: 05:30
- L2 hf ramp burst: 06:00, 17:10
- H1 hf ramp burst: 17:15
- BWS 3.6m/s error block
1.5392 - 17:29 (0.1pa 110) - elec noise: H1, H2, H4, L2
- wnrs noise: L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 100.000 100.000 100.000 100.000 x 99.363 98.862 99.408 98.422   PR - 105177 Daily log
- Uptimes: 252-253
- Reprocess: 252-253
- All: 252-253

Trip return
- Unload truck
- Update bwd pr
- Confirm VTP & BWD are ok
- Write trip reports
- Update pfo group

Michael photo request
- Search all photos, unable to find the one he is looking for
- Respond to his email

Lab safety
- Respond to dennis with suggested times

L1 spiky bdf data
- Investigate high frequency spiky data recently found on L1 bdf channel
- Find a correlation between the appearance of this and the BWD channel offset
- Confirm that spiky data stopped when BWD was fixed yesterday
9/12/2014 255 - fwdr time issue - H4 BDA 20s spikes
- H4 ramp burst: 03:50
- H1 ramp burst: 07:00
- H2 hf ramp burst: 05:45
- L2 hf ramp burst: 07:25
- H1 hf ramp burst: 17:05
- H4 hf ramp burst: 22:35
- BWS 3.6m/s error block
1.8179 - elec noise: H4
- wnrs noise: L1, L2
- 8hz bump: L1-L4
- all clk diff too large
- 8 frames from I57 to IDC missing
- 8 frames from I57 to NDC missing
- 8 frames from I57 to UCSD missing
  98.815 100.000 99.815 99.815 x 99.901 100.000 99.985 99.981 x 99.361 98.866 99.410 98.427   Daily log
- all: 254
9/13/2014 256 - H4 BDA 20s spikes
- H4 ramp burst: 04:50
- H1 ramp burst: 07:50
- H4 hf ramp burst: 01:00, 23:40
- H2 hf ramp burst: 06:40
- L2 hf ramp burst: 08:25
- H1 hf ramp burst: 16:50
- BWS 3.6m/s error block
2.162 - elec noise: H4
- wnrs noise: L2, L3
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.909 100.000 99.986 99.983 x 99.363 98.871 99.412 98.433  
9/14/2014 257 - H4 BDA 20s spikes
- H4 ramp burst: 05:20
- H4 hf ramp burst: 00:50, 03:10, 21:40
- H1 hf ramp burst: 08:00, 16:30
- H2 hf ramp burst: 07:00, 17:15
- L2 hf ramp burst: 08:45, 15:55
- BWS 3.6m/s error block
1.9332 - elec noise: H1, H4
- wnrs noise: L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.915 100.000 99.987 99.985 x 99.366 98.875 99.414 98.439  
9/15/2014 258 - H4 BDA 20s spikes
- H4 ramp burst: 04:35
- H1 ramp burst: 07:30
- H4 hf ramp burst: 02:20, 21:30, 23:25
- H2 hf ramp burst: 07:00, 17:15
- L2 hf ramp burst: 10:05, 15:45
- H1 hf ramp burst: 16:45
- BWS 3.6m/s error block
2.1571 - 21:02 (0.2pa 229) - elec noise: H4
- wnrs noise: L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.921 100.000 99.988 99.986 x 99.368 98.880 99.417 98.446   Daily log
- All: 255-257

Safety
- Ask dennis if we could reschedule

Michael
– status meeting

Jay
- Wants part break down for 18m rosette
- Talking about building a 9th site
- Talked about wilan radios
- He will check on batteries – should have been delivered
- Vault for 9th site

I57 upgades
- Email jay with some discussion topics on the wnrs
- Email jay with reminder to check on the batteries
- Email jay with history and suggested radio vendors
- Email jay with digitizer quote question and mention getting a 6-channel
- Email jay asking about 56 vault performance
- Email shawn asking for pv quote
- Inform jay that I have requested a quote for that same 06’ upgrade system
9/16/2014 259 - H4 BDA 20s spikes
- H4 ramp burst: 06:15
- H1 ramp burst: 09:30
- H2 hf ramp burst: 01:00, 08:40, 17:35
- H4 hf ramp burst: 03:35
- L2 hf ramp burst: 11:10, 16:00
- H1 hf ramp burst: 17:00
- BWS 3.6m/s error block
1.4373 - elec noise: H2, H4, L2
- wnrs noise: L1, L2, H1
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.926 100.000 99.988 99.987 x 99.371 98.884 99.419 98.452  
9/17/2014 260 - H4 BDA 20s spikes
- H4 ramp burst: 02:35
- H1 ramp burst: 07:00
- H4 ramp burst: 22:25
- L2 ramp burst: 22:40
- H2 hf ramp burst: 06:40, 17:30
- L2 hf ramp burst: 08:35, 16:15
- H1 hf ramp burst: 17:20
- BWS 3.6m/s error block
1.2525 - elec noise: H1, H2, H4, L2
- wnrs noise: L2
- 8hz bump: L1-L4
- all clk diff too large   100.000 100.000 100.000 100.000 x 99.930 100.000 99.989 99.988 x 99.373 98.888 99.421 98.458   Daily log
- All: 258-259

truck
- Give michael lease rates on truck from motor pool and suggest we not fix our truck roof

9th element
- Ask jay what he thought about running afar’s with wilans

Safety inspection
9/18/2014 261 - H4 BDA 20s spikes
- H4 hf ramp burst: 03:00
- H1 hf ramp burst: 07:50
- H2 hf ramp burst: 07:40
- L2 hf ramp burst: 08:20
1.5435 - elec noise: H1-H2, L2
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 2 frames showing H1 GPS receiver unlocked
  100.000 100.000 100.000 100.000 x 99.934 100.000 99.990 99.988 x 99.375 98.892 99.423 98.464   Daily log
- All: 260
9/19/2014 262 - H4 BDA 20s spikes
- H1 hf ramp burst: 16:40
- H2 hf ramp burst: 16:30
2.2891 - elec noise:
- wnrs noise: H1-H3, L1-L3
- 8hz bump: L1-L4
- all except H2, H3 clk diff too large
- 3 frames showing H1 GPS receiver unlocked
  100.000 100.000 100.000 100.000 x 99.938 100.000 99.990 99.989 x 99.378 98.897 99.426 98.469   Prep for trip
- Michael Training
- Fuel Truck
- Clean lab
9/20/2014 263 - H4 BDA 20s spikes
- L2 single spike: 14:10
- L1 hf ramp burst: 03:00
- L2 hf ramp burst: 04:35
- H1 hf ramp burst: 06:20
- H2 hf ramp burst: 06:20
1.3052 - elec noise: L2
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 2 frames showing H1 GPS receiver unlocked
  100.000 100.000 100.000 100.000 x 99.941 100.000 99.991 99.990 x 99.380 98.901 99.428 98.475  
9/21/2014 264 - H4 BDA 20s spikes
- L1 single spike: 00:10
- H1 hf ramp burst: 05:40, 18:00
- H2 hf ramp burst: 06:55, 17:30
1.499 - elec noise:
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- all clk diff too large
- 3 frames showing H1 GPS receiver unlocked
  100.000 100.000 100.000 100.000 x 99.944 100.000 99.991 99.990 x 99.382 98.905 99.430 98.481  
9/22/2014 265 - Umoya rcvr offline due to full disk - H4 BDA 20s spikes
- L2 hf ramp burst: 04:45
- H1 hf ramp burst: 05:30, 18:20
- H2 hf ramp burst: 06:10
- L3 hf ramp burst: 11:10
1.3725 - elec noise:
- wnrs noise: H1, L1-L2
- 8hz bump: L1-L4
- N/A   100.000 100.000 100.000 100.000 x 99.946 100.000 99.992 99.991 x 99.385 98.909 99.432 98.487  
9/23/2014 266 - Umoya rcvr restarted - Vandenberg launch: 14:45 - H4 BDA 20s spikes
- H1 hf ramp burst: 05:40
- L2 hf ramp burst: 05:30
- H2 hf ramp burst: 06:30
1.4786 - elec noise:
- wnrs noise: L1-L2
- 8hz bump: L1-L4
- all clk diff too large
- 2 frames showing H1 GPS receiver unlocked
- 1090 frames from I57 to UCSD backfilled
- 1090 frames from profile2 to UCSD backfilled
- 526 frames from profile2 to UCSD duplicated
  100.000 100.000 100.000 100.000 x 99.948 100.000 99.992 99.991 x 99.387 98.913 99.434 98.493  
9/24/2014 267 - H4 BDA 20s spikes
- H4 ramp burst: 02:45
- H1 ramp burst: 06:30
- H4 hf ramp burst: 00:45, 01:40, 23:00
- H1 hf ramp burst: 06:30, 17:10
- H2 hf ramp burst: 08:00
- L2 hf ramp burst: 06:30
1.6348 - 18:04 (0.1pa 109) - elec noise: H1, H4
- wnrs noise: L1-L2
- 8hz bump: L1-L4
- all clk diff too large
- 2 frames showing H1 GPS receiver unlocked
  100.000 100.000 100.000 100.000 x 99.951 100.000 99.992 99.992 x 99.389 98.917 99.436 98.498  
9/25/2014 268 - Umoya rcvr offline due to full disk - H4 BDA 20s spikes
- H1 ramp burst: 08:00
- H4 hf ramp burst: 00;40, 02:20, 03:15, 22:30
- L2 hf ramp burst: 08:50
- H2 hf ramp burst: 13:20
1.5176 - elec noise: H1, H2, H4, L2
- wnrs noise: L1-L3
- 8hz bump: L1-l4
- N/A   100.000 100.000 100.000 100.000 x 99.953 100.000 99.993 99.992 x 99.392 98.921 99.438 98.504  
9/26/2014 269 - Umoya rcvr offline due to full disk - H4 BDA 20s spikes
- H4 hf ramp burst: 00:20, 03:20
1.7071 - elec noise: H1, H4, L2
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- N/A   100.000 100.000 100.000 100.000 x 99.954 100.000 99.993 99.992 x 99.394 98.925 99.441 98.510  
9/27/2014 270 - 07:30 hpwren problems
- Umoya rcvr offline due to full disk
- H4 BDA 20s spikes
- H4 hf ramp burst: 02:10
2.0943 - elec noise: H1, H4, L2
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- N/A   100.000 100.000 100.000 100.000 x 99.956 100.000 99.993 99.993 x 99.396 98.929 99.443 98.515  
9/28/2014 271 - Umoya rcvr offline due to full disk - H4 BDA 20s spikes 1.9482 - elec noise: L2
- wnrs noise: L1-L3
- 8hz bump: L1-L4
- N/A   99.861 100.000 99.861 99.861 x 99.953 100.000 99.988 99.988 x 99.398 98.933 99.444 98.520  
9/29/2014 272 - Umoya full disk repaired and computer restarted - H4 BDA 20s spikes
- L2 ramp burst: 13:10, 15:00
- H1 hf ramp burst: 02:40, 20;50
- H2 hf ramp burst: 23:40
1.2091 - elec noise: L2
- wnrs noise: H1, L1-L2
- 8hz bump: L1-L4
- N/A   100.000 100.000 100.000 100.000 x 99.954 100.000 99.989 99.988 x 99.400 98.937 99.446 98.526   Umoya receiver problems
- Discovered that the source of the problem was low disk space.
- Moved all 2013 waveforms and frame data to second partitition (local)
- Reboot computer
- Waited for I57 and profile2 to backfill

Daily log
- Uptimes: 261-271
- Reprocess: 269-271 (modify to not operate on logs due to umoya backfill)

UAF
- Respond to Jay about cable length
9/30/2014 273 - fwdr time issue - H4 BDA 20s spikes
- H2 hf ramp burst: 02:00
- H1 hf ramp burst: 03:20
1.1654 - 00:33 (0.1pa 226) - elec noise: H2
- wnrs noise: H1, L1-L2
- 8hz bump: L1-l4
- all clk diff too large
- 3 frames showing H1 GPS receiver unlocked
- 2 frames from I57 to IDC delayed
- 4 frames from I57 to IDC duplicated
- 4 frames from I57 to IDC missing
- 1 frame from I57 to NDC delayed
- 4 frames from I57 to NDC duplicated
- 4 frames from I57 to NDC missing
- 1 frame from I57 to UCSD delayed
- 4 frames from I57 to UCSD duplicated
- 4 frmes from I57 to UCSD missing
  99.815 100.000 99.815 99.815 x 99.950 100.000 99.983 99.982 x 99.402 98.941 99.448 98.530   Daily log
- Uptimes: 272

I57 contract
- Tell michael I’m ready for upgrades call, discuss briefly later on phone

Batteries
- Receive 6, check tracking and confirm this is complete order, ask jay if he ordered 24
                    x x x x x x x x x x x x x x